### v2

# Application of the Lovász-Schrijver Lift-and-Project Operator to Compact Stable Set Integer Programs

Abstract
The Lovász theta function $\theta(G)$ provides a very good upper bound on the stability number of a graph $G$. It can be computed in polynomial time by solving a semidefinite program (SDP), which also turns out to be fairly tractable in practice. Consequently, $\theta(G)$ achieves a hard-to-beat trade-off between computational effort and strength of the bound. Indeed, several attempts to improve the theta bound are documented, mainly based on playing around the application of the $N_{+}(\cdot)$ lifting operator of Lovász and Schrijver to the classical formulation of the maximum stable set problem. Experience shows that solving such SDP-s often struggles against practical intractability and requires highly specialized methods. We investigate the application of such an operator to two different linear formulations based on clique and nodal inequalities, respectively. Fewer inequalities describe these two and yet guarantee that the resulting SDP bound is at least as strong as $\theta(G)$. Our computational experience, including larger graphs than those previously documented, shows that upper bounds stronger than $\theta(G)$ can be accessed by a reasonable additional effort using the clique-based formulation on sparse graphs and the nodal-based one on dense graphs.
Keywords semidefinite programming $\cdot$ lift-and-project operator $\cdot$ stable set problem;
MSC 90C22 $\cdot$ 90C27 $\cdot$ 05C69;
1 Introduction
Given a simple undirected graph $G=(V,E)$, with vertex set $V$ and edge set $E$, a stable (or independent) set in $G$ is a set of pairwise non-adjacent vertices. A fundamental problem in combinatorial optimization is the Stable Set Problem (SSP), which consists of computing a stable set of maximum cardinality [33, 50]. This value, denoted by $\alpha(G)$, is the stability number of $G$. A weighted version of this problem can be considered if a weight vector $w\in\mathbb{Q}_{+}^{n}$ is given. SSP is NP-hard in the strong sense and even hard to approximate [35], it is equivalent to the Maximum Clique and Set Packing problems [53] and has a wide range of applications in practical contexts [10, 11, 16, 14, 57, 68]. SSP is naturally formulated as a binary program by the so-called edge inequalities:
$\alpha_{w}(G)=\max$ $\sum_{i\in V}w_{i}x_{i}$
s.t. $\;\;x_{i}+x_{j}\leq 1,\;$$\{i,j\}\in E,$
$x_{i}\in\{0,1\},$ $i\in V.$
As is customary in the literature, we denote by
[…]
the fractional stable set polytope, while
[…]
denotes the stable set polytope, that is, the convex hull of the incidence vectors of all stable sets in $G$. The intense study of STAB($G$) has led, since the early seventies, to the identification of several classes of valid inequalities, often associated with specific “facet-producing” sub-graphs [33, 11, 2, 15]. In practice, however, linear relaxations based on these inequalities often yielded rather weak upper bounds on the stability number of unstructured graphs. As a consequence, basic branch-and-cut algorithms are not particularly effective [49] and much more sophisticated cut generation procedures [30, 31, 20, 19, 56, 55] are required to reduce the performance gap with fast combinatorial exact algorithms, such as those devised in [58, 52, 62].
In his seminal paper, Lovász [42] introduced the celebrated theta number of a graph, denoted by $\theta(G)$. It represents an upper bound for $\alpha(G)$, which can be computed in polynomial time (up to arbitrary precision) by solving a Semidefinite Program (SDP) [33]. Since its introduction, $\theta(G)$, along with the associated non-polyhedral convex theta body $\operatorname{TH}(G)$ [33], turned out to be powerful tools. For instance, they allowed to prove that the SSP can be solved in polynomial time when $G$ is a perfect graph [33]. Indeed, the practical relevance of $\theta(G)$ is no less. On one side, it often turns out to be a significantly stronger upper bound than those from linear relaxations [28, 39]. On the other side, the structure of the associated SDP relaxation often mitigates numerical difficulties that typically affect SDP algorithms. As a consequence, $\theta(G)$ realizes a very good trade-off between the quality of the upper bound and computational effort, which makes it attractive to be incorporated within branch-and-bound algorithms, as studied by, e.g., Wilson [67] and more recently by Gaar et al. [25].
In a successive landmark paper Lovász and Schrijver [43] introduced the $N_{+}(\cdot)$ lift-and-project operator, which can be applied to the LP relaxation of any 0-1 LP and returns a stronger SDP relaxation. Its application to $\operatorname{FRAC}(G)$ has also been intensely investigated and has resulted tighter than the Lovász theta relaxation. From a theoretical perspective, its strength has been documented by showing implications of (additional) facet defining inequalities of the stable set polytope [43, 27]. Moreover, as noted in [7, 8], it provides an effective tool to identify other classes of graphs for which the SSP can be solved in polynomial time. Recently, graph classes yielding stable set polytopes with high rank with respect to $N_{+}(\cdot)$ have been studied in [1].
From the computational point of view, things are somewhat more involved. Experiments conducted with $N_{+}(\operatorname{FRAC}(G))$ [21, 12] and related variants [23, 34] show that these relaxations may yield stronger upper bounds than $\theta(G)$ but at the price of considerably increasing the computational effort. In fact, solving these large-scale SDPs requires specialized methods and large instances can hardly be managed. Other methods, based on strengthening SDP relaxations by linear inequalities not related to the $N_{+}(\cdot)$ operator, have also been presented by Gaar and Rendl [24] and Locatelli [41]. These achieve strong bounds but still require a significant computational effort.
Overall, we currently face a dichotomic picture, where even minimal progress from the Lovász $\theta$ bound may be paid with a very high computational burden. This paper aims to present new SDP relaxations that can bridge such a gap. These are derived by applying the $N_{+}(\cdot)$ operator to alternative LP relaxations. The first relaxation is obtained by replacing edge inequalities with (clearly stronger) clique inequalities. This is unconventional since the linear description may have exponentially many inequalities, making the approach somehow unattractive from theoretical and practical perspectives. However, we show that a natural selection of clique inequalities allows for more compact, tractable, yet stronger, relaxations. An opposite rationale derives the other SDP relaxations: accepting to start from more compact but weaker LP relaxations and relying on the ability of the $N_{+}(\cdot)$ operator to recover their potential weakness. Based on this rationale, we investigate a hierarchy of SDP relaxations where the primary level consists of applying the operator to a surrogate relaxation of $\operatorname{FRAC}(G)$, introduced in Della Croce and Tadei [22], with $O(|V|)$ (so-called nodal) inequalities. Coefficient strengthening procedures then derive two progressively stronger relaxations, the first of which takes polynomial time while the second requires the solution of NP-hard (sub-)problems to be built.
In this work, we first prove that all the proposed SDP relaxations are at least as strong as the Lovász theta relaxation and draw a complete theoretical picture by comparing them to each other and to $N_{+}(\operatorname{FRAC}(G))$. Then, we discuss implementation issues in handling them and illustrate extensive experiments where the practical strength of relaxations is evaluated and contrasted to the theoretical expectation. We show that significant progress with respect to $\theta(G)$ is achieved in several cases. Relaxations obtained from clique inequalities are more effective for sparse graphs, while those from nodal inequalities are the best as graph density grows. Furthermore, the increase of computational workload required to improve the $\theta$ bound is significantly smoother than that of previous approaches. Indeed, this increased efficiency allowed us experimentation based on significantly larger graphs than previously documented. This experience also sheds new light on the practical potential of the $N_{+}(\cdot)$ operator, which emerges more clearly than in the studies documented so far. All the material and code presented is available at https://github.com/febattista/SDP_lift_and_project and organized to facilitate the reproduction of relaxations and experiments.
2 Preliminaries
This paper will adopt the following standard notation. The set of integers $\{1,...,k\}$ is denoted by $[k]$. The complement of a graph $G=(V,E)$ is the graph ${\bar{G}}=\big{(}V,{\bar{E}}\big{)}$, where ${\bar{E}}=\big{\{}\{i,j\}\subset V:\,\{i,j\}\in(V\times V)\setminus E\big{\}}$. Given a vertex $i\in V$, we let $\Gamma_{G}(i)$ denote the set of neighbours of $i$ in $G$, i.e., $\Gamma_{G}(i)=\big{\{}j\in V\setminus\{i\}:\{i,j\}\in E\big{\}}$. When no confusion arises, we let $\Gamma_{G}(i)=\Gamma(i)$. Finally, given a vertex $S\subseteq V$, we let $G[S]=\big{(}S,E(S)\big{)}$ denote the subgraph of $G$ induced by $S$, and $r(S)$ denote its stability number $\alpha\big{(}G[S]\big{)}$ (the so-called rank of $S$). We also let ${\cal S}_{n}$ be the set of real symmetric square matrices of order $n$ and ${\cal S}^{+}_{n}~{}\subset~{}{\cal S}_{n}$ denotes the cone of those which are positive semidefinite (PSD). PSDness of a matrix $Y\in{\cal S}^{+}_{n}$ is also denoted by $Y\succeq 0$. Given any square matrix $X\in\mathbb{R}^{n\times n}$ we denote by $(X)_{ij}$ the element at the $i$-th row and the $j$-th column, by $\operatorname{diag}(X)\in\mathbb{R}^{n}$ its main diagonal and by $\operatorname{rank}(X)$ its rank. We denote by $\left\langle X,Y\right\rangle=\operatorname{tr}(XY)$ the standard inner product in ${\cal S}_{n}$, where $\operatorname{tr}(A)$ is the trace of the square matrix $A$. At last, we denote with $J$ the all-ones square matrix and $e$ the all-ones vector; the dimensions should be clear from the context where not explicit.
2.1 The lift-and-project operator of Lovász and Schrijver
We now review the general statement of the $N_{+}(\cdot)$ operator to make the presentation self-consistent, referring the reader to [18, 21, 32, 43] for an exhaustive treatment.
Let us consider a 0-1 linear program of the form
$\text{max}\left\{c^{\top}x\ |\ Ax\leq b,\ x_{i}\in\{0,1\}\text{ for }i\in I=\{1,\ldots,n\}\right\},$ (IP)
where $c\in\mathbb{R}^{n}$, the inequality system $Ax\leq b$ has $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$ and contains the inequalities $0\leq x_{i}\leq 1$, for $i\in[n]$. Each linear inequality of the system $Ax\leq b$ is denoted by $(a^{k})^{\top}x\leq b^{k}$, for $k\in[m]$. The polytope $P=\{x\in\mathbb{R}^{n}\ |\ Ax\leq b\}$ is the linear relaxation of (IP) and $P_{I}=\operatorname{conv}\{x\in\{0,1\}^{n}\ |\ Ax\leq b\}$ denotes the convex hull of the feasible solutions to (IP).
The application of the lift-and-project operator to $P$ involves the three following steps:
Step 1 (Lifting)
Generate the quadratic inequalities
$x_{i}((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m],$ (2)
$(1-x_{i})((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m].$ (3)
which are valid for $P$ as $x_{i}\geq 0$ and $1-x_{i}\geq 0$ for all $x\in P$. These inequalities can be linearized through the introduction of the matrix variable $X=xx^{\top}$ and rewriting (2) and (3) as:
$\sum_{j\in I}a^{k}_{j}X_{ij}-b^{k}x_{i}$ $\leq 0\quad i\in[n],k\in[m],$ (4)
$\sum_{j\in I}a^{k}_{j}x_{j}-b^{k}-\sum_{j\in I}a^{k}_{j}X_{ij}+b^{k}x_{i}$ $\leq 0\quad i\in[n],k\in[m],$ (5)
$X$ $=xx^{\top}.$
Then, the quadratic constraint $X=xx^{\top}$ is relaxed with $X-xx^{\top}\succeq 0$. By applying the Schur’s complement definition $X-xx^{\top}\succeq 0$ is equivalent to
$\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\succeq 0.$ (6)
Step 2 (Strengthening)
Add to constraints (4)–(6)
$X_{ii}=x_{i}\qquad i\in[n],$ (7)
as any binary vector $x\in P_{I}$ satisfies $x_{i}=x_{i}^{2}$, whereas such equality is not guaranteed for points in $P$. This yields to the definition of the convex set
[…]
where $M_{+}(\cdot)$ is referred here as lift operator.
Step 3 (Projection)
The projection of $M_{+}(P)$ onto the original $x$-space is
${N_{+}(P)}=\left\{x\in\mathbb{R}^{n}\ \Big{|}\ \exists\ Y\in M_{+}(\text{P})\text{ with }\begin{pmatrix}1\\ x\end{pmatrix}=\text{diag}(Y)\right\}.$
[…]
Remark 1.
Let $(a^{k})^{\top}x\leq b^{k}$, for $k\in[m]$, be an inequality in the linear system defining $P$, then from the linear inequalities defining $M_{+}(P)$ one can sum (4) and (5), for any $i\in[n]$, to recover the original constraint.
Moreover, Bodur et al. [9] noted that the definition of $M_{+}(P)$ and $N_{+}(P)$ does not depend on the representation of $P$ but only on the set of points in $P$. Hence, the following property holds:
Property 1.
Given two polytopes $P,P^{\prime}$ with $P^{\prime}\subseteq P$ then ${M_{+}(P^{\prime})}\subseteq{M_{+}(P)}$
Additionally, given any polytope $P$ containing $\operatorname{STAB}(G)$, Lovász and Schrijver [43] provided necessary conditions for an inequality $a^{\top}x\leq b$ to be valid for $N_{+}(P)$.
Lemma 1.
Let $a^{\top}x\leq b$ be a valid inequality for $\operatorname{STAB}(G)$. W.l.o.g. assume $a\in\mathbb{R}^{n}_{+}$ and $b\in\mathbb{R}_{+}$. If $a^{\top}x\leq b$ is valid for $P\cap\{x:x_{i}=1\}$ for all $i\in V$, then $a^{\top}x\leq b$ is valid for $N_{+}(P)$.
3 Continuous relaxations of the SSP
In this section we recall known results about linear and semidefinite relaxation of the SSP. These are widely researched and we will only concentrate on elements relevant to our study.
3.1 Linear relaxations of the SSP
Several classes of valid inequalities for $\operatorname{STAB}(G)$ have been identified since early seventies [33, 11]. Among them, clique inequalities $\sum_{i\in C}x_{i}\leq 1$, where $C\subseteq V$ is a maximal clique (set of pairwise adjacent vertices), induce facets of STAB$(G)$ [53]. As customary in the literature, we define the polytope
${\operatorname{QSTAB}(G)}=\{x\in[0,1]^{|V|}:\sum_{i\in C}x_{i}\leq 1,\;C\in{\cal K}\}$
where ${\cal K}$ denotes the collection of all maximal cliques of $G$. As clique inequalities clearly imply edge inequalities, STAB$(G)\subseteq$ $\operatorname{QSTAB}(G)$subseteq FRAC$(G)$ and inclusion is generally strict [33]. Clique inequalities in $G$ can be exponential in number, and the associated separation problem is NP-hard in the strong sense. However, effective separation heuristics make this class of inequalities computationally easy to handle, as shown in [45].
The 0-1 linear program
$\max$ $\sum\limits_{i\in V}w_{i}x_{i}$
s.t. $\sum\limits_{i\in C}x_{i}\leq\,1\quad(C\in{\mathcal{C}})$ (8)
$x_{i}\in\{0,1\}\quad(i\in V).$
is a valid formulation for the SSP if ${\mathcal{C}}\subseteq{\cal K}$ is any collection of maximal cliques covering all the edges of $G$ (i.e., both the endpoints of every edge are contained in at least one clique of $\mathcal{C}$). Extending the standard notation, we denote by $\operatorname{QSTAB}(G,\mathcal{C})$ the polytope associated with its continuous relaxation. It’s not hard to see that ${\operatorname{QSTAB}(G)}\subseteq{\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC(G)}$. $\operatorname{QSTAB}(G,\mathcal{C})$ has a number of clique inequalities bounded by $|E|$ and it typically contains considerably fewer inequalities than $\operatorname{FRAC(G)}$. In fact, formulations based on a greedily computed minimum size collection $\mathcal{C}$ are competitive in practice, as shown in several studies (see, e.g., Letchford et al. [40] and references therein).
Other rank inequalities of the form $\sum_{i\in S}x_{i}\leq\alpha(G[S])$ have been studied for vertex subsets $S\subseteq V$ inducing special graphs (see e.g., Borndörfer [11] and Giandomenico et al. [28] for a survey). A well-known example arises when $S=H$, $|H|\geq 5$, induces a chordless cycle of odd cardinality, yielding the odd-hole inequality $\sum_{i\in H}x_{i}\leq\lfloor\frac{|H|}{2}\rfloor$. Among other subgraphs, we recall webs and antiwebs introduced in [63]. Let $p$ and $q$ be integers satisfying $p>2q+1$ and $q>1$, and use arithmetic modulo $p$. A $(p,q)$--web is a graph with vertex set $\{1,...,p\}$ and with edges from $i$ to $\{i+q,\ldots,i-q\}$, for every $1\leq i\leq p$. A $(p,q)$--antiweb is the complement of a $(p,q)$--web. The web inequalities take the form $\sum_{i\in W}x_{i}\leq q$ for every vertex set $W$ inducing a $(p,q)$-web, and the antiweb inequalities take the form $\sum_{i\in AW}x_{i}\lfloor p/q\rfloor$ for every vertex set $AW$ inducing a $(p,q)$-antiweb.
A different modeling rationale has been pursued by Della Croce and Tadei [22] who considered the surrogate relaxation of formulation (1) obtained by summing up, for every $i\in V$, the edge inequalities over all $j\in\Gamma(i)$. This yields the so called nodal inequality $\sum_{j\in\Gamma(i)}x_{j}+|\Gamma(i)|x_{i}\,\leq\,|\Gamma(i)|$. Murray and Church [48] observed that this relaxation can be strengthened by replacing $|\Gamma(i)|$ with any value $r_{i}\geq\alpha(G[\Gamma(i)])$, which returns a class of 0-1 linear programs of the form
$\max$ $\,\,\sum\limits_{i\in V}w_{i}x_{i}$
s.t. $\sum\limits_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i}\quad(i\in V)$ (9)
$x_{i}\in\{0,1\}\quad(i\in V).$
The strongest formulation, corresponding to $r_{i}=\alpha(G[\Gamma(i)])$, has been extensively investigated in [40]. Of course, computing coefficients $r_{i}=\alpha(G[\Gamma(i)])$ is NP-hard, as it amounts to solve the SSP on the subgraphs $G[\Gamma(i)]$. Nevertheless, Letchford et al. [40] showed that these formulations can be handled efficiently for many graphs of interest. Here, we are interested in a further choice of coefficients $r_{i}$, namely, $r_{i}=\theta(G[\Gamma(i)])$ where $\theta(G[\Gamma(i)])$ is the Lovász [42] $\theta$ number of the graph $G[\Gamma(i)]$. This represents an upper bound to $\alpha(G[\Gamma(i)])$ which can be computed in polynomial time up to an arbitrary precision, as reviewed in Section 3.2.1. Then, we introduce the following notation
${\operatorname{NOD}(G,r)}=\left\{x\in[0,1]^{|V|}:\sum_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i},\hskip 5.0pti\in V\right\}$
to denote the polytope defined from (9) (and non-negativity), where $r\in\{\Gamma,\theta,\alpha\}$ specifies the coefficient used. Clearly, we have ${\operatorname{NOD}(G,\alpha)}\subseteq{\operatorname{NOD}(G,\theta)}\subseteq{\operatorname{NOD}(G,\Gamma)}$. It is worth noting that nodal inequalities in general do not imply edge inequalities, even in the case $r=\alpha$. On the other hand, if $r=\alpha$, they imply the wheel inequalities (see Borndörfer [11] for their definition) which are not implied by clique inequalities.
3.2 SDP relaxations of the SSP
Three well-known SDP relaxations are related to our study.
3.2.1 The Lovász theta relaxation
The Lovász theta relaxation has been introduced in 1979, yet we refer to the presentation of [43]. Introduce the quadratic variable $x_{ij}$, representing the product $x_{i}x_{j}$ for all $\{i,j\}\in V\times V$ and let $X=xx^{\top}$ be the associated symmetric matrix of order $n=|V|$. An upper bound to $\alpha(G)$ is given by
$\theta(G,w)=\max$ $\;\sum_{i\in V}w_{i}x_{i}$ (SDP-$\theta$)
s.t. $\;X_{ii}=x_{i},\quad i\in V,$
$\;X_{ij}=0,\quad\{i,j\}\in E,$
$\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\succeq 0.$ (10)
We simply denote it by $\theta(G)$ in the unweighted case. The projection $\operatorname{TH}(G)$ of the feasible region of (SDP-$\theta$) onto the subspace of the $x_{i}$ variables is known as the theta body, a convex but not polyhedral set in general. Remarkably, Grötschel et al. [33] proved that $\operatorname{STAB}(G)\subseteq\operatorname{TH}(G)\subseteq{\operatorname{QSTAB}(G)}$, where equality holds if and only if $G$ is perfect. This implies that there exists a polynomial-time separation algorithm for a class of inequalities which includes all clique inequalities. A thorough comparison between the original formulation in [42] and (SDP-$\theta$) is illustrated in [26]. $\theta(G)$ can be computed in polynomial time up to an arbitrary precision and often provides a strong bound to the stability number $\alpha(G)$, typically better than those obtained from linear relaxations see, e.g., [38, 69, 28, 30]. Another interesting feature of (SDP-$\theta$) concerns with its computational behavior. In fact, (SDP-$\theta$) is handled better than similarly sized unstructured SDPs by general SDP solvers. Moreover, several algorithms have been designed to solve it efficiently as in, e.g., [54, 44, 29]. Therefore, $\theta(G)$ achieves a good compromise between strength of the upper bound and computational burden. From a theoretical viewpoint, Busygin and Pasechnik [13] showed that, unless P=NP, no polynomially computable upper bound to $\alpha(G)$ which is provably smaller than $\theta(G)$ can be found. In this context, several stronger relaxations have been obtained by adding linear inequalities to (SDP-$\theta$), as reviewed in the following subsections.
3.2.2 The Schrijver relaxation
Schrijver [59] observed that the PSD condition (10) does not imply the non-negativity on $X$. Hence, by adding the inequalities
$X_{ij}\geq 0,\quad\forall\{i,j\}\in\bar{E},$ (SDP-$\theta^{+}$)
to (SDP-$\theta$), one yields a model that we denote as (SDP-$\theta^{+}$) and the corresponding upper bound as $\theta^{+}(G)$. Accordingly, $\operatorname{TH}^{+}(G)$ denotes the convex set obtained by its projection onto the subspace of $x_{i}$ variables.
3.2.3 The Lovász-Schrijver relaxation
The Lovász-Schrijver relaxation for the SSP is obtained by applying the operator $N_{+}(\cdot)$ to $\operatorname{FRAC}(G)$. $N_{+}(\operatorname{FRAC}(G))$ corresponds to (SDP-$\theta^{+}$) plus the following linear inequalities:
$X_{ij}$ $\geq 0,$ $\{i,j\}\in\bar{E},$ (11)
$X_{ik}+X_{jk}$ $\leq x_{k},$$\{i,j\}\in E,k\neq i,j,$ (12)
$x_{i}+x_{j}+x_{k}$ $\leq 1+X_{ik}+X_{jk},$$\{i,j\}\in E,k\neq i,j.$ (13)
Lovász and Schrijver [43] assessed the theoretical strength of $N_{+}(\operatorname{FRAC}(G))$ by showing that it satisfies several well-known classes of valid inequalities for $\operatorname{STAB}(G)$. Namely, they proved that $N_{+}(\operatorname{FRAC}(G))$ satisfies all clique, odd cycle, odd antihole and odd wheel inequalities. Giandomenico and Letchford [27] showed that, in addition, it satisfies all web inequalities. Optimizing $\sum_{i\in V}x_{i}$ over $N_{+}(\operatorname{FRAC}(G))$ returns an upper bound to $\alpha(G)$ that we denote as $\lambda(G)$. From a more general perspective, it is interesting to observe the key role of the positive semidefiniteness constraint. In fact, Burer and Vandenbussche [12] compared $\lambda(G)$ with the optimal value of the corresponding LP relaxation obtained by the Sherali-Adams procedure, observing a substantial improvement due to it.
Handling such a strong relaxation in practice is not straightforward. The first related computational results, obtained by a lift-and-project cutting plane method, appear in [3]. Results on optimising directly over $M_{+}(\operatorname{FRAC}(G))$ are illustrated in [21] and, more recently, in [12] using a specialised augmented Lagrangian method. The resulting upper bounds to $\alpha(G)$ may be considerably better than $\theta^{+}(G)$ but at the expense of a large increase in running times. In practice, this relaxation turns out to be hardly tractable for graphs of medium/large size.
3.2.4 Further SDP relaxations
Variants of the previous relaxations are presented in [34, 23]. In the former, extra triangle inequalities are added to $N_{+}(\operatorname{FRAC}(G))$ to obtain a very strong relaxation. In the latter, relaxations are specialised with respect to graph density to improve efficiency. We refer to [26] for an exhaustive review of these methods. Two SDP-related methodologies are also presented in the literature. In [24] strong SDP relaxations of several combinatorial optimization problems are investigated by exploiting the exact subgraph inequalities. Despite the power of this technique, the experience reveals that strengthening the natural SDP relaxation of the SSP is typically harder than other problems, e.g., max-cut, and confirms that improving the $\theta$ bound towards $\alpha(G)$, even by a small amount, is actually challenging. Another effective method (presented in the context of the clique number) based on adding non-valid inequalities has been developed in [41]. This turns out to achieve quite strong bounds for both structured and random graphs with up to 150 vertices.
Summarizing, the evidence from state-of-the-art methodologies is that achieving even a small improvement of $\theta(G)$ requires a considerable additional computational cost. Namely, adding extra linear inequalities in the SDP models yields much harder SDPs and does require specialized methods. In the next section we introduce new relaxations with the goal of bridging such a gap.
4 New SDP relaxations
We now apply the lift operator $M_{+}(\cdot)$ to $\operatorname{QSTAB}(G,\mathcal{C})$ and $\operatorname{NOD}(G,r)$ and compare the resulting SDP relaxations. Let us first observe that all of them contain variable bound constraints, the lifting step of which consists in multiplying $0\leq x_{i}\leq 1$ by $x_{j}$ and $(1-x_{j})$, for $i,j\in V,i<j$:
$x_{i}x_{j}\rightarrow$ $\quad X_{ij}\geq 0,$ (14)
$x_{i}(1-x_{j})\rightarrow$ $\quad x_{i}\geq X_{ij},$ (15)
$(1-x_{i})x_{j}\rightarrow$ $\quad x_{j}\geq X_{ij},$ (16)
$(1-x_{i})(1-x_{j})\rightarrow$ $\quad x_{i}+x_{j}-X_{ij}\leq 1.$ (17)
The linear inequalities (14)–(17) are known as McCormick [46] inequalities and describe the convex hull of the set $\{(x_{i},x_{j},y_{ij})\in\{0,1\}^{3}:y_{ij}=x_{i}x_{j}\}$. It is easy to see that they are not implied by the PSD condition (6) and should be explicitly added to the relaxations illustrated below unless implied by other constraints.
4.1 Relaxation $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$
In order to investigate $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ we first describe the application of the operator $M_{+}(\cdot)$ to $\operatorname{QSTAB}(G)$. The lifting step consists in multiplying each clique inequality $\sum_{i\in C}x_{i}\leq 1$, $C\in{\cal K}$, by $x_{j}$ and $(1-x_{j})$, for all $j\in V$. Then, adding the condition $X_{ii}=x_{i}$, one obtains:
$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C\setminus\{j\}}X_{ij}\leq 0,$ if $j\in C$, (18)
$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C}X_{ij}-x_{j}\leq 0,$ if $j\not\in C$, (19)
$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}x_{i}-\sum\limits_{i\in C\setminus\{j\}}X_{ij}\leq 1,$ if $j\in C$, (20)
$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}\leq 1,$ if $j\not\in C$. (21)
Inequalities (19), called clique-variable inequalities, have been introduced in [28]. One can observe that
  1. 1.
    Constraints (18) imply $X_{ij}=0$ for all $\{i,j\}\in E$ and constraints (20) reduce to the corresponding clique inequality $\sum_{i\in C}x_{i}\leq 1,C\in{\cal K}$, which is already implied from (19) and (21) by Remark 1;
  2. 2.
    $X_{ij}=0$ for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (19) imply inequalities (15) and (16);
  3. 3.
    Constraints (18), (19) and (21) imply (17).
Then, the description of $M_{+}(\operatorname{QSTAB}(G))$ reduces to
$\begin{split}{M_{+}(\operatorname{QSTAB}(G))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ \sum\limits_{i\in C}X_{ij}-x_{j}&\leq 0,\quad C\in\mathcal{K},j\in V\setminus C,\\ \sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}&\leq 1,\quad C\in\mathcal{K},j\in V\setminus C,\\ x_{i}&=X_{ii},\quad i\in V,\\ X_{ij}&=0,\quad\{i,j\}\in E,\\ X_{ij}&\geq 0,\quad\{i,j\}\in{\bar{E}}\\ x_{i}&\geq 0,\quad i\in V\bigg{\}}.\end{split}$ (22)
Moreover, one has:
Theorem 1.
${M_{+}(\operatorname{QSTAB}(G))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$.
Proof Since ${\operatorname{QSTAB}(G)}\subseteq\operatorname{FRAC}(G)$, the statement follows from Property 1. ∎
We denote by $\mu(G)=\max\left\{\sum_{i\in V}x_{i}:\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\in{M_{+}(\operatorname{QSTAB}(G))}\right\}$ the resulting upper bound to $\alpha(G)$. The following example shows that the containment ${M_{+}(\operatorname{QSTAB}(G))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$ can be strict.
Example 1.
Let $G$ be the $(10,3)$-antiweb graph of Figure 2(a). We have $\mu(G)=3$. Differently, optimizing the cardinality objective function $\sum_{i\in V}x_{i}$ over $N_{+}(\operatorname{FRAC}(G))$ one obtains the optimal solution $x_{F}=(0.3106,\ldots,0.3106)$ with value $z^{*}_{F}=3.106$. Therefore, $x_{F}\in{N_{+}(\operatorname{FRAC}(G))}$ lies outside $N_{+}(\operatorname{QSTAB}(G))$.
The example indeed suggests a general result. In fact, we can show that, unlike $N_{+}(\operatorname{FRAC}(G))$, relaxation $N_{+}(\operatorname{QSTAB}(G))$ implies the class of antiweb inequalities.
Theorem 2.
$N_{+}(\operatorname{QSTAB}(G))$ implies antiweb inequalities
Proof Let $G$ be the $(p,q)$-antiweb graph with $V=\left\{1,\ldots,p\right\}$. In what follows, we use arithmetic modulo $p$. Note that, by definition, any consecutive $q$ nodes in $G$ form a clique. By denoting $\alpha=\left\lfloor\frac{p}{q}\right\rfloor$, the $(p,q)$-antiweb inequality is
$\sum\limits_{i\in V}x_{i}\leq\alpha.$
Given any $i\in V$, we denote with $G\setminus i$ the subgraph induced by the vertex set
$V(G\setminus i)=V\setminus(\left\{i\right\}\cup\Gamma(i))=\left\{i+q,\ i+q+1,\dots,\ i-q-1,\ i-q\right\}.$
The goal is to prove that inequality
$\sum\limits_{j\in V(G\setminus i)}x_{j}\leq\alpha-1,$ (23)
is valid for ${\operatorname{QSTAB}(G\setminus i)}$, for all $i\in V$, then the theorem’s statement will follow from Lemma 1. Let $i$ be any node in $V$ and consider the following $\alpha-1$ cliques in $G\setminus i$:
$C^{1}$ =$=\left\{i+q,\ i+q+1,\dots,\ i+2q-1\right\},$
$C^{2}$ =$=\left\{i+2q,\ i+2q+1,\dots,\ i+3q-1\right\},$
$\vdots$
$C^{k}$ =$=\left\{i+kq,\ i+kq+1,\dots,i+(k+1)q-1\right\},$
$\vdots$
$C^{\alpha-1}$ =$=\left\{i+(\alpha-1)q,\ i+(\alpha-1)q+1,\dots,\ i-q\right\}.$
Note that the set $\left\{C^{1},\dots,C^{\alpha-1}\right\}$ forms a partition of $V(G\setminus i)$, i.e., $\bigcup_{k=1,\dots,\alpha-1}C^{k}=V(G\setminus i)$ and $C^{\ell}\cap C^{k}=\emptyset$, for all $\ell,k=1,\dots,\alpha-1$, with $\ell\not=k$. Hence, the sum of clique inequalities $\sum_{j\in C^{k}}x_{j}\leq 1$, for $k=1,\dots,\alpha-1$ yield the inequality
$\sum\limits_{i+q\leq j\leq i-q}x_{j}=\sum\limits_{j\in V(G\setminus i)}x_{j}\leq\alpha-1,$
which is valid for $\operatorname{QSTAB}(G\setminus i)$. By symmetry, this argument is independent of the choice of $i$. ∎
Theorem 1 completes the theoretical picture for $N_{+}(\operatorname{QSTAB}(G))$ along the line drawn in [43]. In fact, all of the traditional combinatorial inequalities, namely, clique, odd cycle, odd antihole, odd wheel, web and antiweb inequalities turn out to be satisfied by $N_{+}(\operatorname{QSTAB}(G))$. It is interesting to mention that Giandomenico et al. [28] proved the same result for a non-compact linear relaxation obtained by the application of the $N(K,K)$ lifting operator by Lovász and Schrijver to $\operatorname{QSTAB}(G)$. The comparison between $N(K,K)$ and $N_{+}(\cdot)$ was indeed suggested by Lovász and Schrijver [43] as an interesting open issue.
Another insight of the progress of $N_{+}(\operatorname{QSTAB}(G))$ towards $N_{+}(\operatorname{FRAC}(G))$ stems from the analysis of graph $\hat{G}$ of Figure 1 which has been shown in [6] that $N_{+}(\operatorname{FRAC}(\hat{G}))$ gives rise to a non-polyhedral relaxation. Here we have $\lambda(\hat{G})=2.146$. Now, optimizing over $N_{+}(\operatorname{QSTAB}(\hat{G}))$ one gets $\mu(\hat{G})=2$. This implies that $N_{+}(\operatorname{QSTAB}(\hat{G}))$ is actually polyhedral.
Figure 1: The $\hat{G}$ graph.
From a computational perspective, a major issue of $N_{+}(\operatorname{QSTAB}(G))$ is the linear description of $\operatorname{QSTAB}(G)$ that, in general, contains exponentially many inequalities. However, as ${\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$, we still have ${M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$ and one can consider the convex set:
[…]
Note that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ still implies McCormick inequalities (14)–(17). Remarkably, it can be built in polynomial time as a suitable set $\cal C$ can be determined by a greedy algorithm (e.g., the one proposed in Letchford et al. [40]), and the number of inequalities describing $\operatorname{QSTAB}(G,\mathcal{C})$ is bounded by $|E|$ (for increasing graph density, it is also considerably smaller than $|E|$). Optimizing $\sum_{i\in V}x_{i}$ over $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ returns an upper bound that we denote as $\mu(G,{\cal C})$. We will experiment with such a relaxation in Section 6.
4.2 Relaxation $N_{+}(\operatorname{NOD}(G,r))$
Let us now consider a generic nodal inequality, with $r_{i}\geq\alpha(G[\Gamma(i)])$, for all $i\in V$. The lifting step along with $x_{i}=X_{ii}$ returns:
$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}X_{ih}\leq 0,$ if $j=i$ (24)
$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $(1-r_{i})x_{j}+\sum_{h\in\Gamma(i)\setminus j}X_{jh}+r_{i}X_{ij}\leq 0,$ if $j\in\Gamma(i)$ (25)
$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}X_{jh}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,$ if $j\not\in\Gamma(i)$ (26)
$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}(x_{h}-X_{ih})+r_{i}x_{i}\leq r_{i},$ if $j=i$ (27)
$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)\setminus j}(x_{h}-X_{jh})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\in\Gamma(i)$ (28)
$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}(x_{h}-X_{jh})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\not\in\Gamma(i)$ (29)
One can observe that:
  1. 1.
    Constraints (24) imply $X_{ij}=0$ for all $\{i,j\}\in E$; this is a relevant fact if one recalls that edge inequalities are not implied by nodal inequalities (9);
  2. 2.
    $X_{ij}=0$, for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (26) imply inequalities (15) and (16);
  3. 3.
    $X_{ij}=0$, for all $\{i,j\}\in E$, inequalities (15) and (16) and constraints (28) imply edge inequalities (and (17) for $\{i,j\}\in E$);
  4. 4.
    By $X_{ij}=0$, for all $\{i,j\}\in E$, (27) boils down to the original nodal inequality (9) and can be dropped, since Remark 1 can be applied to both pairs (25), (28) and/or (26), (29);
  5. 5.
    $X_{ij}\geq 0$, for all $\{i,j\}\in\bar{E}$, inequalities (15) and (16) and constraints (29) imply (17) for $\{i,j\}\in\bar{E}$.
Thanks to these properties $M_{+}(\operatorname{NOD}(G,r))$ boils down to
$\begin{split}{M_{+}(\operatorname{NOD}(G,r))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ (1-r_{i})x_{j}+\sum_{h\in\Gamma(i)\setminus j}X_{jh}\leq 0,&\quad i\in V,j\in\Gamma(i),\\ \sum_{h\in\Gamma(i)}X_{jh}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,&\quad i\in V,j\not\in\Gamma(i),\\ \sum_{h\in\Gamma(i)\setminus j}(x_{h}-X_{hj})+r_{i}x_{i}+r_{i}x_{j}\leq r_{i},&\quad i\in V,j\in\Gamma(i),\\ \sum_{h\in\Gamma(i)}(x_{h}-X_{hj})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},&\quad i\in V,j\not\in\Gamma(i),\\ x_{i}=X_{ii},&\quad i\in V,\\ X_{ij}=0,&\quad\{i,j\}\in E,\\ X_{ij}\geq 0,&\quad\{i,j\}\in{\bar{E}},\\ x_{i}\geq 0,&\quad i\in V\bigg{\}}.\end{split}$ (30)
Then, $M_{+}(\operatorname{NOD}(G,r))$ can be obtained by adding linear inequalities (25), (26), (28), (29) and $X_{ij}\geq 0$, for $\{i,j\}\in\bar{E}$, to (SDP-$\theta$) (it is straightforward to observe that inequalities (25) are implied by Mc-Cormick inequalities for $r_{i}=\Gamma$). Therefore, its projection $N_{+}(\operatorname{NOD}(G,r))$ onto the $x$-space is at least as strong as the Schrijver relaxation:
Theorem 3.
${N_{+}(\operatorname{NOD}(G,r))}\subseteq\operatorname{TH}^{+}(G)$.
Which also implies:
Corollary 1.
$N_{+}(\operatorname{NOD}(G,r))$ implies all clique inequalities.
Corollary 2.
$\nu(G,r)=\max\left\{\sum_{i\in V}x_{i}:\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\in{M_{+}(\operatorname{NOD}(G,r))}\right\}\leq\theta^{+}(G)$.
Theorem 3 holds in particular for $r_{i}=|\Gamma(i)|$ corresponding to the case where the initial formulation is the surrogate relaxation of $\operatorname{FRAC(G)}$. Even in this case, applying the $N_{+}(\cdot)$ operator can recover the initial formulation’s weakness and yield an upper bound that is at least as good as Schrijver’s bound. Furthermore, a hierarchy of relaxations derives from strengthening the lifting coefficient $r_{i}$, that is, looking at $r_{i}=\theta(G[\Gamma(i)])$ and $r_{i}=\alpha(G[\Gamma(i)])$. Thanks to Property 1 we have
Theorem 4.
${M_{+}(\operatorname{NOD}(G,\alpha))}\subseteq{M_{+}(\operatorname{NOD}(G,\theta))}\subseteq{M_{+}(\operatorname{NOD}(G,\Gamma))}$.
This hierarchy poses a complexity issue. The strongest relaxation comes from applying the operator to an LP relaxation, the construction of which, in general, cannot be carried out in polynomial time. An interesting polynomial special case arises when the subgraphs induced by neighbor sets are perfect and, according to the result in [33], in this case, we have $\theta(G[\Gamma(i)])=\alpha(G[\Gamma(i)])$, for all $i\in V$. Differently, ${M_{+}(\operatorname{NOD}(G,\theta))}$ can always be constructed (and solved) in polynomial time. Section 6 documents the practical trade-off between strength and computational tractability for all relaxations in this hierarchy.
4.3 Comparison among relaxations
A natural question concerns with relationships between $N_{+}(\operatorname{FRAC}(G))$, $N_{+}(\operatorname{QSTAB}(G))$ and $N_{+}(\operatorname{NOD}(G,r))$. The following example shows that neither $N_{+}(\operatorname{NOD}(G,r))$ contains $N_{+}(\operatorname{FRAC}(G))$ nor the reverse.
Example 2.
Consider the odd-hole with seven vertices of Figure (2(b)). According to [43], $N_{+}(\operatorname{FRAC}(G))$ satisfies the facet-defining inequality $\sum_{i=1}^{8}x_{i}\leq 3$ which implies $\lambda(G)=3$. Differently, the optimal value from $N_{+}(\operatorname{NOD}(G,r))$ is $\nu(G,\theta)=\nu(G,\alpha)=3.317$ corresponding to solution $(0.47395,\ldots,0.47395)$, which therefore falls outside $N_{+}(\operatorname{FRAC}(G))$. However, one can find a feasible solution to $N_{+}(\operatorname{FRAC}(G))$ which is outside $N_{+}(\operatorname{NOD}(G,r))$ as well. Consider the $(10,3)$-antiweb graph $G$ of Figure (2(a)). Here, $\nu(G,\theta)=\nu(G,\alpha)=\alpha(G)=3$. On the contrary, optimizing over $N_{+}(\operatorname{FRAC}(G))$ we obtain solution
$x_{F}=(0.31055,\dots,0,31055)$
with value $3.105$ which is then outside $N_{+}(\operatorname{NOD}(G,r))$, for $r\in\{\theta,\alpha\}$.
11112222333344445555666677778888999910101010
(a) $(10,3)$–antiweb
1111222233334444555566667777
(b) 7-hole
Figure 2: Two example graphs.
This example also shows that the odd-hole inequalities are not implied by $N_{+}(\operatorname{NOD}(G,r))$ even with $r=\alpha$. This appears as a weakness indicator of $N_{+}(\operatorname{NOD}(G,r))$ if one recalls that odd-hole inequalities are contained in the first Chvàtal closure of the edge formulation 1 [53]. The overall theoretical picture is summarized in Figure 3, and the above results seem to suggest that $N_{+}(\operatorname{QSTAB}(G))$ (as well as $N_{+}(\operatorname{FRAC}(G))$) should be stronger than $N_{+}(\operatorname{NOD}(G,r))$. Interestingly, the computational analysis will reveal significant supremacy of the latter as graphs get dense, for $r\in\{\theta,\alpha\}$.
Symbol Relaxation
$\lambda(G)$ $N_{+}(\operatorname{FRAC}(G))$
$\mu(G)$ $N_{+}(\operatorname{QSTAB}(G))$
$\mu(G,{\cal C})$ $N_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$
$\nu(G,r)$ $N_{+}(\operatorname{NOD}(G,r))$, for $r\in\{\Gamma,\theta,\alpha\}$
Table 1: Notation for SDPs optimal values.
Figure 3: Containment relationships among relaxations.
5 Implementation
The software implemented for this work, along with the datasets used in the experiments (see Section 6), is available at https://github.com/febattista/SDP_lift_and_project. The repository comprises the following Python (.py) and MATLAB (.m) files:
  • LinearFormulations.py: this module creates all linear formulations presented in Section 3.1 and exports them in standard .lp format. It relies on the max-clique solver cliquer [52] and ADMMsolver.py (see below) to compute coefficients for nodal inequalities (9).
  • SDPLifting.py: starting from a linear formulation in standard .lp format, this module implements the $M_{+}(\cdot)$ operator described in Section 2.1, exporting the resulting SDP in the .mat MATLAB format.
  • ADMMsolver.py: this module implements the SDP solver proposed in [4, 5] and is exclusively employed as a subroutine during the LP formulation.
  • kelley_cutting_plane.m: this module implements the cutting plane method described in Algorithm 1.
The following Python scripts are provided to simplify the use of the software:
  • install.py downloads and builds all required dependencies (i.e. SDPNAL+ [61] and cliquer [52]).
  • parameters.py sets up the computational experiments, allowing for (i) selection of the datasets to be used and (ii) tuning of all the parameters of Algorithm 1 and the SDP solvers involved.
  • analyze_results.py collects the data from the experiments and replicates all the tables presented in this work.
Selection of the SDP solver
[…]
Model building
[…]
Graph
brock800_1
brock800_2
brock800_3
brock800_4
C500-9
p_hat300-1
p_hat300-2
p_hat300-3
p_hat500-1
p_hat500-2
p_hat500-3
p_hat700-1
p_hat700-2
p_hat700-3
sanr400_0.5
sanr400_0.7
$\theta_{i}$ Min 2.94 2.83 3.07 2.81 0.76 1.96 1.39 0.44 4.63 2.61 0.58 8.89 5.30 0.77 1.27 0.69
Max 4.98 4.97 4.94 4.91 10.92 7.43 20.35 9.33 10.73 78.22 13.28 25.81 123.65 24.58 2.64 2.04
Average 3.93 3.90 3.93 3.93 2.40 3.53 6.14 1.86 7.66 15.95 3.84 15.06 31.25 6.98 1.78 1.09
$\alpha_{i}$ Min 0.89 1.08 0.68 0.63 $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ 0.01 $<0.01$
Max 28.87 34.41 26.17 25.91 $<0.01$ $<0.01$ 0.02 0.02 0.02 1.83 2.74 0.07 62.41 108.48 0.04 0.13
Average 6.00 6.17 5.93 6.04 $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ 0.10 0.10 0.02 1.64 2.78 0.02 0.03
Table 2: CPU time statistics for computing coefficients $\theta$ and $\alpha$ for (9) on DIMACS graphs.
Lifting operation
The function m_plus_lifting() in SDPLifting.py carries out Steps 1 and 2 of the procedure outlined in Section 2.1. Initially, it parses the polytope $P$, provided as input through a standard .lp format text file. Then, it construct the symmetric matrices corresponding to constraints (4), (5) and (7). Finally, the model is exported in .mat file format, ready to be loaded into MATLAB and solved using SDPNAL+.
SDP optimization
In general, we provide as input to SDPNAL+ SDPs of the form
$\operatorname{opt}(\operatorname{R})=\max_{Y\in\operatorname{R}}\langle C,Y\rangle,$
where
$\operatorname{R}=\left\{Y\in{\cal S}^{+}_{n+1}\ :\ Y\geq 0,\langle A_{i},Y\rangle\leq b_{i},\ i\in[m]\right\},$
and $m\in\mathbb{N}$, $C,A_{1},\ldots,A_{m}\in{\cal S}_{n+1}$, $b\in\mathbb{R}^{m}$. In a preliminary experiment we observed that SDPNAL+ converges quickly on (SDP-$\theta^{+}$) while it requires very large (and practically unacceptable) convergence times on $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ relaxations. This is due to the number of linear constraints in the former, and their density in the latter. However, we proved that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ are contained in (SDP-$\theta^{+}$), thus one can resort to a classical cutting plane method to optimize over them. The cutting plane algorithm starts optimizing over (SDP-$\theta^{+}$). At the $k$-th iteration, given the optimal solution $Y^{k}\in{\cal S}^{+}_{n+1}$, one can compute
$V_{i}(Y^{k})=\langle A_{i},Y^{k}\rangle-b_{i},\ i\in[m].$
[…]
Algorithm 1 Cutting-plane scheme
[…]
2:Initialize $\Pi^{0}\leftarrow{\cal S}_{n+1},\Delta^{0}=+\infty,\ k=0$
[…]
4:while $Y^{k}$ violates constraints in $\operatorname{R}$ or $\Delta^{k}>\delta$ do
5:     Compute the vector $V(Y^{k})$ of violations and $I_{c}(Y^{k})$
6:     $\Pi^{k+1}\leftarrow\Pi^{k}\cap\left\{Y\in{\cal S}_{n+1}\ :\ \langle A_{i},Y\rangle\leq b_{i},\ i\in I_{c}(Y^{k})\right\}$
[…]
8:     $\Delta^{k+1}\leftarrow|p^{k+1}-p^{k}|$
9:     $k\leftarrow k+1$
10:end while
6 Computational Experiments
We now illustrate the results obtained by solving relaxations $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ for $r\in\{\Gamma,\theta,\alpha\}$ with Algorithm 1. Experiments are designed to address three main questions:
  • (i)
    Can relaxations based on operator $M_{+}(\cdot)$ provide upper bounds for $\alpha(G)$ that are significantly stronger than $\theta(G)$?
  • (ii)
    What is the additional computational cost for achieving such an improvement?
  • (iii)
    How much does the choice of $r$ affect the quality of the bound from $M_{+}(\operatorname{NOD}(G,r))$?
As a byproduct, we document the quality of the bound provided by $M_{+}(\operatorname{NOD}(G,\theta))$, that is, the strongest bound which can be computed in polynomial time in our hierarchy.
Experiments are run on a computer with Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz, 256GB RAM, under OS Ubuntu 16.04.7 LTS (GNU/Linux 4.15.0-128-generic x86_64). After preliminary tuning experiments, the parameters of Algorithm 1 have been set to $\epsilon=10^{-3},\delta=10^{-1}$ and $c=1000$ with 7200 seconds time limit (excluding formulation building time). The precision of the SDP solver SDPNAL+ is set to $10^{-6}$. The precision of ADMMsolver.py for evaluating coefficients $\theta_{i}$ is set to $10^{-4}$. Denoting by $\operatorname{opt}(\operatorname{R})$ the optimal (with the given precision) value of relaxation $\operatorname{R}$ computed by Algorithm 1 at termination, tables below report, for each relaxation, four statistics: the percentage gap to $\alpha(G)$, that is, $\frac{\operatorname{opt}(\operatorname{R})-\alpha(G)}{\alpha(G)}\%$; the number of cutting plane iterations and the corresponding CPU time; the number of added cuts. The latter is detailed for each class of inequalities and a class is not reported if none of its members has ever been detected as violated.
6.1 Instances
The numerical experiments are based on the following two collections of graphs, available at https://github.com/febattista/SDP_lift_and_project in the standard edge-list format.
Random graphs
A Erdös–Rényi graph $G(n,p)$ has $n$ vertices and each edge appears with probability $p\in[0,1]$. We have considered the collection of graphs from [40], generated by
$n\in\{200,250,300\},\quad p\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}.$
where, for each combination of $n$ and $p$, five instances with different random seeds have been created for a total of 135 graphs.
DIMACS graphs
Graphs from the Second DIMACS Implementation Challenge [37]. These form the standard benchmark for max-clique algorithms. We complemented the graphs to convert the max-clique instances into SSP instances (DSJC* graphs are not complemented, as these belong to the “coloring” benchmark set). We excluded the c-fat, johnson, and san graphs because the integrality gap is completely closed by linear relaxations, so they do not provide information to our analysis. We also dropped keller5 and the largest and p-hat instances (with 1000 and 1500 vertices) as the SDP relaxations turned out to be too large to be handled. The final collection includes 38 graphs. Statistics about the time required to compute lifting coefficients of linear formulations $N_{+}(\operatorname{NOD}(G,\theta))$ and $N_{+}(\operatorname{NOD}(G,\alpha))$ are reported in Table 2. Recall that their computation can be carried out in parallel, so the impact on the overall time is often negligible.
6.2 Experiment 1: random graphs
Table 3 collects gap statistics for Erdös–Rényi graphs, where the first column reports the integrality gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$. Tables 4 and 5 report cutting plane iterations, CPU times, and, respectively, the number of added cuts for each class of inequalities. Let us first notice that $\lambda(G)\simeq\theta^{+}(G)$ in all cases but $n=200$, $p\in\{0.1,0.9\}$, evidencing a rather disappointing behavior of $M_{+}(\operatorname{FRAC}(G))$. On the contrary, replacing edge inequalities with a compact collection of clique inequalities in the initial formulation pays off, as $\mu(G,{\cal C})$ is always smaller than $\lambda(G)$. The average portion of additional gap closed by $\mu(G,{\cal C})$ w.r.t. $\lambda(G)$ is $6.7\%$. This is achieved by a few cutting plane iterations, adding about one to three thousand inequalities of class (19), resulting in an average $80\%$ CPU time increase. Interestingly, no inequalities in the class (21) have been detected as violated.
This experiment clearly classifies relaxations strength as a function of edge probability $p$ and graph size $n$. When $p\leq 0.4$, $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ turns out to be the strongest relaxation and the sparser the graph, the larger the gap closed by $\mu(G,{C})$ w.r.t. other relaxations, while formulations based on nodal inequalities fail to improve on $\theta^{+}(G)$. However, the advantage of $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ over other relaxations tends to thin out as $n$ increases. The outcome is completely reversed when $p\geq 0.5$, where relaxations based on nodal inequalities are more competitive and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$ is the clear winner. $\nu(G,\alpha)$ closes $45.3\%$ of the gap left by $\lambda(G)$ and $40.0\%$ of the one left by $\mu(G,{C})$, on average. To our knowledge, such an improvement on $\theta^{+}(G)$ has rarely been documented in the literature on SDP approaches, even for smaller graphs. Another nice evidence is that this advantage of $M_{+}(\operatorname{NOD}(G,\alpha))$ is independent of the graph size. Observe that, also in this case, one class of inequalities is predominant, namely, (25). Notice also that some violated inequalities in (28) and (29) have only been detected for the densest graphs $p=0.9$. Very good upper bounds are also achieved by $M_{+}(\operatorname{NOD}(G,\theta))$ when $p\geq 0.7$. In fact, $\nu(G,\theta)$ remarkably improves on $\theta^{+}(G)$ and, unlike $\nu(G,\alpha)$, can be computed in polynomial time. Also, the computational effort required to achieve strong bounds is reasonable. Differently, the weakest relaxation $M_{+}(\operatorname{NOD}(G,\Gamma))$ always returns $\nu(G,\Gamma)=\theta^{+}(G)$, that is, never closes additional gap in our tests. Finally, notice that for highly dense graphs, the gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$ may be smaller than that of the weakest SDP relaxations. However, it is largely improved by the strongest SDP relaxation, which confirms the operator’s power and the PSD constraint’s key role.
Graph % gap
$n$ $p$ $\operatorname{NOD}(G,\alpha)$ $\theta^{+}(G)$ $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
200 0.1 46.957 19.330 18.723 16.033 19.330 19.330 19.330
0.2 102.539 27.924 27.912 26.060 27.924 27.924 27.924
0.3 93.334 30.774 30.773 29.393 30.774 30.774 30.774
0.4 73.543 30.550 30.550 29.322 30.550 30.550 30.537
0.5 56.527 30.597 30.597 29.355 30.597 30.593 29.540
0.6 39.889 25.708 25.708 24.318 25.708 25.143 20.138
0.7 30.543 25.579 25.579 23.679 25.579 19.897 12.494
0.8 9.720 13.474 13.474 10.819 13.474 2.871 1.836
0.9 9.600 10.123 10.120 5.346 10.123 5.057 5.057
250 0.1 64.521 27.371 27.112 24.908 27.371 27.371 27.371
0.2 118.348 35.462 35.459 34.284 35.462 35.462 35.462
0.3 103.026 39.136 39.136 38.370 39.136 39.136 39.136
0.4 78.520 38.335 38.335 37.737 38.335 38.335 38.312
0.5 54.333 33.563 33.563 32.943 33.563 33.563 31.421
0.6 46.178 39.416 39.416 38.669 39.416 39.163 28.532
0.7 39.114 38.531 38.531 37.459 38.531 36.305 23.679
0.8 19.867 20.012 20.012 18.473 20.012 9.419 4.433
0.9 5.950 11.717 11.716 9.524 11.717 3.219 3.219
300 0.1 75.661 31.124 31.047 29.389 31.124 31.124 31.124
0.2 137.919 46.238 46.237 45.551 46.238 46.238 46.238
0.3 113.088 48.426 48.426 48.033 48.426 48.426 48.426
0.4 89.040 50.228 50.228 49.918 50.228 50.228 50.166
0.5 62.433 46.061 46.061 45.791 46.061 46.061 41.439
0.6 42.224 39.956 39.956 39.595 39.956 39.848 26.444
0.7 23.650 31.619 31.619 31.059 31.619 28.272 11.347
0.8 23.333 29.809 29.809 28.718 29.809 19.807 11.667
0.9 7.090 23.160 23.160 20.318 23.160 5.512 5.486
Table 3: $N_{+}(\cdot)$ bounds on random instances. The % gap of the LP relaxation ${\operatorname{NOD}(G,\alpha)}$ is reported as a reference. The best gaps are shown in boldface.
Graph $\theta^{+}(G)$ $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
$n$ $p$ Time # iter Time # iter Time # iter Time # iter Time # iter Time
200 0.1 7.24 1.6 26.64 3.2 48.24 0.2 8.74 0.2 8.64 0.2 8.76
0.2 4.29 1.0 9.10 2.0 13.96 0.0 4.34 0.0 4.34 0.0 4.34
0.3 2.64 0.8 4.43 2.0 8.53 0.0 2.71 0.0 2.71 0.0 2.71
0.4 1.74 0.2 2.65 2.0 6.51 0.0 1.82 0.2 2.13 1.0 3.77
0.5 1.64 0.0 2.46 2.0 6.37 0.0 1.74 1.0 3.60 1.0 12.12
0.6 1.82 0.0 2.85 2.0 7.00 0.0 1.93 1.0 6.56 2.0 22.02
0.7 2.18 0.0 3.26 2.0 10.90 0.0 2.30 2.0 29.37 2.4 58.03
0.8 3.05 0.0 4.42 2.0 14.14 0.0 3.18 1.4 49.62 1.4 40.79
0.9 5.43 0.6 10.33 2.0 37.73 0.0 5.57 2.6 236.08 2.6 236.57
250 0.1 9.62 1.0 28.92 3.0 56.76 0.0 9.68 0.0 9.68 0.0 9.68
0.2 5.15 1.0 11.53 2.0 18.05 0.0 5.25 0.0 5.25 0.0 5.25
0.3 4.05 0.4 5.82 2.0 11.34 0.0 4.19 0.0 4.18 0.0 4.18
0.4 2.40 0.2 4.17 1.0 5.12 0.0 2.57 0.0 2.57 1.0 4.95
0.5 2.10 0.0 4.08 1.0 4.78 0.0 2.30 0.6 3.51 2.0 23.23
0.6 2.05 0.0 4.21 1.0 4.79 0.0 2.26 1.0 6.46 3.0 53.63
0.7 2.07 0.0 4.66 1.0 5.87 0.0 2.30 2.0 29.51 3.0 65.98
0.8 2.52 0.0 5.04 1.0 8.39 0.0 2.80 2.8 112.29 2.6 162.95
0.9 5.77 0.2 9.94 1.2 20.54 0.0 6.05 3.0 430.18 3.0 429.29
300 0.1 10.92 1.0 25.16 3.0 68.62 0.0 11.04 0.0 11.03 0.0 11.03
0.2 8.02 0.6 13.87 2.0 25.53 0.0 8.18 0.0 8.19 0.0 8.18
0.3 5.26 0.0 6.88 1.0 9.61 0.0 5.47 0.0 5.47 0.0 5.48
0.4 3.31 0.0 5.72 1.0 6.59 0.0 3.60 0.0 3.58 1.0 7.26
0.5 2.67 0.0 5.62 1.0 5.64 0.0 3.01 0.0 3.00 2.0 31.62
0.6 2.54 0.0 6.24 1.0 5.89 0.0 2.90 1.0 7.51 3.0 66.31
0.7 2.37 0.0 6.86 1.0 5.80 0.0 2.84 2.0 40.36 4.0 135.51
0.8 3.78 0.0 7.42 1.0 10.76 0.0 4.31 2.2 97.62 3.6 197.18
0.9 6.61 0.0 11.98 1.8 30.75 0.0 7.11 4.8 1250.18 5.0 1566.21
Table 4: CPU time and cutting plane’s iterations on random graphs.
Graph Number of violated inequalities added in Algorithm 1
$\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\theta)$ $\nu(G,\alpha)$
$n$ $p$ (12) (19) (25) (26) (28) (29) (25) (26) (28) (29)
200 0.1 1018 3126 0 0 0 0 0 0 0 0
0.2 30 2000 0 0 0 0 0 0 0 0
0.3 2 1517 0 0 0 0 0 0 0 0
0.4 0 1326 0 1 0 0 8 2 0 0
0.5 0 1232 7 1 0 0 569 65 0 0
0.6 0 1213 322 28 0 0 1459 128 0 0
0.7 0 1247 1508 52 0 0 2137 157 0 0
0.8 0 1273 1332 27 0 0 1395 5 0 0
0.9 1 1280 1800 0 567 60 1800 0 567 60
250 0.1 615 3000 0 0 0 0 0 0 0 0
0.2 7 1821 0 0 0 0 0 0 0 0
0.3 0 1281 0 0 0 0 0 0 0 0
0.4 0 972 0 0 0 0 21 3 0 0
0.5 0 907 1 1 0 0 1095 145 0 0
0.6 0 862 201 11 0 0 2799 201 0 0
0.7 0 908 1122 32 0 0 2924 76 0 0
0.8 0 994 2759 41 0 0 2378 122 0 0
0.9 0 755 2400 0 232 24 2400 0 232 24
300 0.1 256 3000 0 0 0 0 0 0 0 0
0.2 1 1462 0 0 0 0 0 0 0 0
0.3 0 946 0 0 0 0 0 0 0 0
0.4 0 657 0 0 0 0 69 8 0 0
0.5 0 534 0 0 0 0 1822 178 0 0
0.6 0 547 128 4 0 0 2828 172 0 0
0.7 0 662 1772 60 0 0 3939 61 0 0
0.8 0 848 2138 62 0 0 3574 26 0 0
0.9 0 1179 4000 0 460 63 4200 0 568 73
Table 5: Number of added cutting planes on random instances.
6.3 Experiment 2: DIMACS instances
The DIMACS test set, a standard for max-clique/stable set studies, includes, for the most part, structured graphs from various applications. The breadth of the graph collection supports the relevance of our analysis. To our knowledge, DIMACS graphs with more than 300-400 vertices have never been documented in SDP-based studies [24, 23, 25, 41]. The graphs we have considered are significantly larger than those in case studies applying the $N_{+}(\cdot)$ operator in [21, 12]. In our experience, experimenting with larger graphs is crucial to draw reliable conclusions about the strength of relaxations for the stable set problem, as it typically tends to degrade with graph size. Looking at Table 6, as already noted for random graphs, $\lambda(G)$ does not prove to be better than $\theta^{+}(G)$: some improvement is observed in two cases (DSJC125-1 and MANN_a27), it is negligible in nine cases and null in the remaining ones. Concerning the relaxations we have proposed, $M_{+}(\operatorname{NOD}(G,\alpha))$ shows the most promising results. $\nu(G,\alpha)$ is the best bound in 11 cases while $\mu(G,{\cal C})$ is the winner in as many as 23 cases. Nevertheless, $\nu(G,\alpha)$ achieves the smallest gap on average, thanks to its remarkable strength on selected instances. According to the evidence gathered from Erdös–Rényi graphs, $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ is more suited for sparse instances, and $M_{+}(\operatorname{NOD}(G,\alpha))$ turns out to be particularly tight as the density increases. $\mu(G,{\cal C})$ is the best bound for the C and MANN collections and the sparsest among p_hat graphs. On the other hand, $\nu(G,\alpha)$ is very strong for DSJC500-5, p_hat300-1, p_hat500-1, p_hat700-1, sanr400_0.5, where it closes more than $50\%$ of the gap left by $\theta^{+}(G)$ and $\lambda(G)$ (about $74\%$ for p_hat-1 graphs). The computational price for this is reported in Table 7: computing $\mu(G,\alpha)$ is about 27 times slower than computing $\theta^{+}(G)$ except p_hat700-1 where it is 185 times slower. But this experiment also shows that the tightness of $M_{+}(\operatorname{NOD}(G,\alpha))$ is less sensitive to the graph size with respect to the other relaxations. This results in an interesting outcome for the hard brock800 collection that, with a density of $35\%$, one expects to be favorable to $M_{+}(\operatorname{FRAC}(G))$ and $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$. $\nu(G,\alpha)$ is the only bound stronger than $\theta^{+}(G)$ on these graphs, and the improvement is significant. Table 7 shows that computing $\nu(G,\alpha)$ is about four times slower than computing $\theta^{+}(G)$ (including the time to build the LP relaxation), an acceptable price for such a valuable result. Table 8 also reveals the key contribution to closing the gap given by inequalities (25).
Also in this experiment, the polynomial bound $\nu(G,\theta)$ improves on $\theta^{+}(G)$ in six cases by a meaningful amount (namely, DSJC125-5, MANN_a27, p_hat300-1, p_hat500-1, p_hat700-1). The comparison between $\nu(G,\theta)$ and $\nu(G,\alpha)$ reveals that strengthening the lifting coefficient pays off, as in eleven cases, $\nu(G,\alpha)$ is significantly smaller than $\nu(G,\theta)$, in other eight cases the improvement is negligible while no improvement is observed in the remaining sixteen cases. On the other hand, $\nu(G,\Gamma)$ has never enhanced $\theta^{+}(G)$ even if a few violated cuts have occasionally been detected. Looking at Tables 7 and 8, one can observe that relevant improvements on $\theta^{+}(G)$ are achieved by adding several inequalities in a few iterations of the cutting plane algorithm. Similarly to what is observed for random graphs, inequalities (21) are only added in a few special graphs, namely the extremely sparse MANN graphs. Differently, classes of inequalities (25) and (26) are relevant to describe $M_{+}(\operatorname{NOD}(G,\theta))$, $M_{+}(\operatorname{NOD}(G,\alpha))$, while no violated members of classes (28) and (29) have ever been found.
Algorithm 1 has shown a nice numerical behavior even for the largest graphs in the DIMACS collection, and re-optimization is cost-effective (Table 7): its extra time is only paid when the cuts yield a significant improvement of the upper bound. In contrast, in the other cases, CPU times do not deviate significantly from those of $\theta^{+}(G)$.
Overall, the computational analysis shows that the new SDP relaxations $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$, may be a viable option to compute upper bounds to $\alpha(G)$ which are remarkably stronger than $\theta(G)$ at a reasonable computational price. Interestingly, some experimental evidence complements the theoretical insights illustrated in Section 4.3, where $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ looks to be likely to achieve the tightest bounds than $M_{+}(\operatorname{NOD}(G,\alpha))$.
7 Conclusions
We have presented a way to improve the classical Lovász $\theta$ bound for the stability number of a graph by applying the Lovász-Schrijver $N_{+}(\cdot)$ lift-and-project operator to tailored LP relaxations. Unlike previous approaches, some of the latter may not be constructed in polynomial time. In the case of clique inequalities, we have to handle exponentially many inequalities (with an associated NP-hard separation problem), while building strong nodal inequalities requires solving the SSP on specific subgraphs. We show that these expedients, if properly handled, are indeed helpful in letting the resulting SDP relaxations improve the $\theta$ bound without excessive extra effort. This study reveals that the Lovász and Schrijver operators can be of practical interest besides being a powerful theoretical tool. Finally, an interesting research direction consists of finding classes of graphs where ${N_{+}(\operatorname{NOD}(G,\theta))}=\operatorname{STAB}(G)$, as for them the SSP would be polynomially solvable.
Acknowledgements
We wish to thank Adam Letchford and Stefano Gualandi for helpful discussions about the preliminary results of this study.
Conflict of interest
The authors declare that they have no conflict of interest.
Graph $\theta^{+}(G)$ $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
Name $d$(%) UB % gap UB % gap UB % gap UB % gap UB % gap UB % gap
brock200_1 25 27.20 29.508 27.20 29.505 27.12 29.119 27.20 29.508 27.20 29.508 27.20 29.508
brock200_2 50 14.13 17.758 14.13 17.758 14.12 17.674 14.13 17.758 14.13 17.757 14.02 16.795
brock200_3 39 18.67 24.479 18.67 24.479 18.65 24.356 18.67 24.479 18.67 24.478 18.67 24.467
brock200_4 34 21.12 24.242 21.12 24.242 21.10 24.101 21.12 24.242 21.12 24.242 21.12 24.242
brock400_1 25 39.33 45.670 39.33 45.670 39.33 45.662 39.33 45.670 39.33 45.670 39.33 45.670
brock400_2 25 39.20 35.160 39.20 35.160 39.20 35.156 39.20 35.160 39.20 35.160 39.20 35.160
brock400_3 25 39.16 26.324 39.16 26.324 39.16 26.310 39.16 26.324 39.16 26.324 39.16 26.324
brock400_4 25 39.23 18.883 39.23 18.883 39.23 18.877 39.23 18.883 39.23 18.883 39.23 18.883
brock800_1 35 41.87 82.032 41.87 82.032 41.87 82.032 41.87 82.032 41.87 82.032 41.55 80.646
brock800_2 35 42.10 75.435 42.10 75.435 42.10 75.435 42.10 75.435 42.10 75.435 41.73 73.896
brock800_3 35 41.88 67.530 41.88 67.530 41.88 67.530 41.88 67.530 41.88 67.530 41.51 66.043
brock800_4 35 42.00 61.541 42.00 61.541 42.00 61.541 42.00 61.541 42.00 61.541 41.61 60.039
C125-9 10 37.55 10.431 36.75 8.074 36.23 6.566 37.55 10.431 37.54 10.424 37.55 10.429
C250-9 10 55.82 26.856 55.71 26.618 55.20 25.446 55.82 26.856 55.82 26.856 55.82 26.856
C500-9 10 83.58 46.630 83.58 46.630 83.51 46.515 83.58 46.630 83.58 46.630 83.58 46.630
DSJC125.1 9 38.04 11.896 37.28 9.650 36.79 8.207 38.04 11.895 38.04 11.881 38.04 11.881
DSJC125.5 50 11.40 14.021 11.40 14.017 11.36 13.598 11.40 14.021 11.38 13.810 11.35 13.531
DSJC125.9 90 4.00 0.000 4.00 0.000 4.00 0.000 4.00 0.000 4.00 0.000 4.00 0.000
DSJC500-5 50 22.57 73.621 22.57 73.621 22.57 73.621 22.57 73.621 22.57 73.621 20.54 58.014
MANN_a9 7 17.48 9.219 17.09 6.811 17.00 6.250 17.48 9.219 17.47 9.201 17.47 9.201
MANN_a27 1 132.76 5.367 131.11 4.057 131.01 3.980 132.76 5.367 131.99 4.751 131.99 4.751
johnson32-2-4 12 16.00 0.000 16.00 0.000 16.00 0.000 16.00 0.000 16.00 0.000 16.00 0.000
keller4 35 13.47 22.417 13.46 22.388 13.39 21.709 13.47 22.417 13.45 22.236 13.45 22.236
p_hat300-1 76 10.02 25.253 10.02 25.253 10.02 25.194 10.02 25.253 9.58 19.691 8.58 7.288
p_hat300-2 51 26.71 6.855 26.58 6.317 26.49 5.950 26.71 6.855 26.70 6.813 26.69 6.768
p_hat300-3 26 40.70 13.057 40.62 12.829 40.44 12.320 40.70 13.057 40.70 13.054 40.70 13.053
p_hat500-1 75 13.01 44.533 13.01 44.533 13.01 44.526 13.01 44.533 12.65 40.583 10.68 18.721
p_hat500-2 50 38.56 48.306 38.44 47.863 38.38 47.598 38.56 48.306 38.56 48.293 38.54 48.225
p_hat500-3 25 57.81 15.622 57.76 15.524 57.67 15.342 57.81 15.622 57.81 15.622 57.81 15.621
p_hat700-1 75 15.05 36.774 15.05 36.774 15.05 36.773 15.05 36.774 14.86 35.047 11.30 2.709
p_hat700-2 50 48.44 10.091 48.30 9.776 48.24 9.646 48.44 10.091 48.44 10.087 48.41 10.023
p_hat700-3 25 71.76 15.734 71.71 15.660 71.64 15.545 71.76 15.734 71.75 15.734 71.75 15.733
sanr200_0.7 30 23.63 31.296 23.63 31.291 23.59 31.048 23.63 31.296 23.63 31.296 23.63 31.296
sanr200_0.9 10 48.90 16.440 48.63 15.786 48.03 14.356 48.90 16.440 48.90 16.440 48.90 16.440
sanr400_0.5 50 20.18 55.217 20.18 55.217 20.18 55.216 20.18 55.217 20.18 55.217 19.10 46.951
sanr400_0.7 30 33.97 61.746 33.97 61.746 33.97 61.741 33.97 61.746 33.97 61.746 33.97 61.746
Means - 30.832 - 30.528 - 30.248 - 30.832 - 30.489 - 27.772
Table 6: Lift-and-project bounds on DIMACS instances. The column $d(\%)$ report the density of each graph.
Graph $\theta^{+}(G)$ $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
Name $d$(%) Time # iter Time # iter Time # iter Time # iter Time # iter Time
brock200_1 25 3.72 1 7.22 1 6.86 0 3.79 0 3.78 0 3.79
brock200_2 50 1.63 0 2.51 1 4.02 0 1.73 1 3.79 1 14.22
brock200_3 39 2.07 1 4.27 1 4.44 0 2.15 1 3.92 1 4.47
brock200_4 34 2.66 0 3.12 1 4.79 0 2.75 0 2.75 1 4.35
brock400_1 25 11.59 0 15.10 1 20.59 0 12.00 0 12.07 0 11.99
brock400_2 25 12.35 0 15.47 1 21.18 0 12.75 0 12.77 0 12.76
brock400_3 25 12.17 0 14.87 1 21.40 0 12.59 0 12.58 0 12.61
brock400_4 25 12.05 0 14.84 1 20.76 0 12.45 0 12.44 0 12.44
brock800_1 35 20.60 0 51.54 0 27.57 0 24.54 0 24.60 2 77.95
brock800_2 35 20.26 0 53.58 0 27.05 0 24.25 0 24.22 2 86.15
brock800_3 35 20.55 0 52.85 0 28.19 0 25.02 0 24.63 2 79.55
brock800_4 35 19.86 0 50.82 0 27.46 0 23.98 0 24.00 2 76.29
C125-9 10 3.46 2 15.68 3 32.44 1 7.10 1 7.44 1 7.55
C250-9 10 10.45 1 32.52 2 42.68 0 10.52 0 10.51 0 10.52
C500-9 10 25.12 1 57.41 1 53.55 0 25.52 0 25.50 0 25.50
DSJC125.1 9 4.38 2 19.28 3 34.02 1 8.38 1 10.27 1 9.02
DSJC125.5 50 1.58 1 3.17 1 3.72 0 1.61 1 4.35 1 4.64
DSJC125.9 90 1.65 0 1.98 0 1.72 0 1.70 0 1.70 0 1.70
DSJC500-5 50 6.35 0 20.49 0 8.85 0 7.71 0 7.83 4 126.12
MANN_a9 7 0.79 1 2.15 1 4.38 0 0.80 1 2.20 1 1.82
MANN_a27 1 9.93 1 73.41 3 1992.72 0 10.00 1 57.22 1 60.32
johnson32-2-4 12 4.16 0 6.92 0 4.28 0 4.58 0 4.58 0 4.60
keller4 35 3.97 1 7.70 1 13.20 0 4.03 1 15.51 1 15.12
p_hat300-1 76 9.07 0 12.62 1 22.93 0 9.49 2 134.81 3 261.52
p_hat300-2 51 80.54 1 220.85 2 221.98 0 80.85 1 300.55 1 279.48
p_hat300-3 26 17.66 1 32.39 2 44.90 0 17.84 1 39.16 1 39.27
p_hat500-1 75 19.08 0 40.72 1 42.03 0 21.13 2 164.90 5 858.88
p_hat500-2 50 201.27 2 661.29 2 652.87 0 202.66 1 580.22 1 893.64
p_hat500-3 25 35.75 1 88.95 2 116.21 0 36.50 1 79.32 1 90.28
p_hat700-1 75 35.53 0 82.10 1 79.12 0 41.34 2 260.41 7 6580.82
p_hat700-2 50 425.90 2 1270.35 2 1409.98 0 429.64 1 920.86 1 1251.71
p_hat700-3 25 81.42 1 204.92 2 278.70 0 83.93 1 189.04 1 204.45
sanr200_0.7 30 3.05 1 5.39 1 5.54 0 3.11 0 3.11 0 3.11
sanr200_0.9 10 7.77 2 29.39 3 43.35 0 7.81 1 15.36 1 14.86
sanr400_0.5 50 4.40 0 11.74 1 8.96 0 5.09 0 5.11 3 77.60
sanr400_0.7 30 7.69 0 11.58 1 12.80 0 8.16 0 8.17 0 8.26
Table 7: CPU time and cutting plane’s iterations on DIMACS graphs.
Number of violated inequalities added in Algorithm 1
Graph $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
Name (12) (19) (21) (26) (25) (26) (25) (26)
brock200_1 10 471 0 0 0 0 0 0
brock200_2 0 81 0 0 5 4 583 76
brock200_3 1 152 0 0 1 1 2 5
brock200_4 0 139 0 0 0 0 1 0
brock400_1 0 34 0 0 0 0 0 0
brock400_2 0 38 0 0 0 0 0 0
brock400_3 0 49 0 0 0 0 0 0
brock400_4 0 33 0 0 0 0 0 0
brock800_1 0 0 0 0 0 0 1845 155
brock800_2 0 0 0 0 0 0 1780 220
brock800_3 0 0 0 0 0 0 1825 175
brock800_4 0 0 0 0 0 0 1840 160
C125-9 1413 2248 3 1 0 35 0 14
C250-9 565 1988 0 0 0 0 0 0
C500-9 1 633 0 0 0 0 0 0
DSJC125.1 1435 2249 11 8 0 36 0 36
DSJC125.5 8 210 0 0 62 17 117 24
DSJC125.9 0 0 0 0 0 0 0 0
DSJC500-5 0 0 0 0 0 0 3772 228
MANN_a9 720 900 36 0 0 504 0 504
MANN_a27 1000 2000 1000 0 0 1000 0 1000
johnson32-2-4 0 0 0 0 0 0 0 0
keller4 48 296 0 0 96 48 96 48
p_hat300-1 0 48 0 0 1911 89 2923 77
p_hat300-2 988 1379 0 0 121 346 329 398
p_hat300-3 761 1878 0 0 0 60 0 75
p_hat500-1 0 13 0 0 1982 18 4959 41
p_hat500-2 1042 1363 0 0 30 288 576 424
p_hat500-3 548 1625 0 0 0 28 0 72
p_hat700-1 0 2 0 0 1960 40 6962 38
p_hat700-2 1107 1549 0 0 16 286 718 282
p_hat700-3 520 1414 0 0 0 47 1 187
sanr200_0.7 3 286 0 0 0 0 0 0
sanr200_0.9 1060 2216 0 0 0 1 0 1
sanr400_0.5 0 3 0 0 0 0 2806 194
sanr400_0.7 0 29 0 0 0 0 0 0
Table 8: Number of added cutting planes on DIMACS instances.

### v3

@@
-Application of the Lovász-Schrijver Lift-and-Project Operator to Compact Stable Set Integer Programs
+Application of the Lovász-Schrijver Operator to Compact Stable Set Integer Programs
 Abstract
@@
 Consequently, $\theta(G)$ achieves a hard-to-beat trade-off between computational effort and strength of the bound.
-Indeed, several attempts to improve the theta bound are documented, mainly based on playing around the application of the $N_{+}(\cdot)$ lifting operator of Lovász and Schrijver to the classical formulation of the maximum stable set problem.
-Experience shows that solving such SDP-s often struggles against practical intractability and requires highly specialized methods.
-We investigate the application of such an operator to two different linear formulations based on clique and nodal inequalities, respectively.
-Fewer inequalities describe these two and yet guarantee that the resulting SDP bound is at least as strong as $\theta(G)$.
+Indeed, several attempts to improve the theta bound are documented, mainly based on playing around the application of the $N_{+}(\cdot)$ lifting operator of Lovász and Schrijver to the classical formulation of the maximum stable set problem (SSP).
+Experience shows that solving such SDPs often struggles against practical intractability and requires highly specialized methods.
+We investigate the application of such an operator to two different linear formulations of the SSP based on clique and nodal inequalities, respectively.
+These two formulations are described by fewer inequalities than the natural formulation based on edge inequalities, yet they guarantee that the resulting SDP bound is at least as strong as $\theta(G)$.
 Our computational experience, including larger graphs than those previously documented, shows that upper bounds stronger than $\theta(G)$ can be accessed by a reasonable additional effort using the clique-based formulation on sparse graphs and the nodal-based one on dense graphs.
@@
 This value, denoted by $\alpha(G)$, is the stability number of $G$.
-A weighted version of this problem can be considered if a weight vector $w\in\mathbb{Q}_{+}^{n}$ is given.
-SSP is NP-hard in the strong sense and even hard to approximate [35], it is equivalent to the Maximum Clique and Set Packing problems [53] and has a wide range of applications in practical contexts [10, 11, 16, 14, 57, 68].
-SSP is naturally formulated as a binary program by the so-called edge inequalities:
-$\alpha_{w}(G)=\max$ $\sum_{i\in V}w_{i}x_{i}$
+For the sake of presentation, we do not explicitly address the weighted version of this problem, where a weight vector $w\in\mathbb{Q}_{+}^{|V|}$ is given, although all the results presented remain valid in that setting.
+The SSP is NP-hard in the strong sense and even hard to approximate [35], it is equivalent to the Maximum Clique and Set Packing problems [53] and has a wide range of applications in different practical contexts, see, e.g., those illustrated in [14, 16, 57] and those surveyed in [10, 68].
+The SSP is naturally formulated as a binary program by the so-called edge inequalities
+$\alpha(G)=\max$ $\sum_{i\in V}x_{i}$
 s.t. $\;\;x_{i}+x_{j}\leq 1,\;$ $\{i,j\}\in E,$
@@
 denotes the stable set polytope, that is, the convex hull of the incidence vectors of all stable sets in $G$.
-The intense study of STAB($G$) has led, since the early seventies, to the identification of several classes of valid inequalities, often associated with specific “facet-producing” sub-graphs [33, 11, 2, 15].
+The intense study of STAB($G$) has led, since the early seventies [2], to the identification of several classes of valid inequalities, often associated with specific “facet-producing” sub-graphs [11, 15, 33].
 In practice, however, linear relaxations based on these inequalities often yielded rather weak upper bounds on the stability number of unstructured graphs.
-As a consequence, basic branch-and-cut algorithms are not particularly effective [49] and much more sophisticated cut generation procedures [30, 31, 20, 19, 56, 55] are required to reduce the performance gap with fast combinatorial exact algorithms, such as those devised in [58, 52, 62].
+As a consequence, basic branch-and-cut algorithms are not particularly effective [49] and much more sophisticated cut generation procedures, see [19, 20, 30, 31, 55, 56], are required to reduce the performance gap with fast combinatorial exact algorithms, such as those devised in [52, 58, 62].
 In his seminal paper, Lovász [42] introduced the celebrated theta number of a graph, denoted by $\theta(G)$.
-It represents an upper bound for $\alpha(G)$, which can be computed in polynomial time (up to arbitrary precision) by solving a Semidefinite Program (SDP) [33].
-Since its introduction, $\theta(G)$, along with the associated non-polyhedral convex theta body $\operatorname{TH}(G)$ [33], turned out to be powerful tools.
-For instance, they allowed to prove that the SSP can be solved in polynomial time when $G$ is a perfect graph [33].
+It represents an upper bound on $\alpha(G)$, which can be computed in polynomial time (up to arbitrary precision) by solving a Semidefinite Program (SDP).
+Since its introduction, $\theta(G)$, along with the associated non-polyhedral convex theta body $\operatorname{TH}(G)$, turned out to be powerful tools.
+For instance, they allowed to prove that the SSP can be solved in polynomial time when $G$ is a perfect graph.
+For the description of these results we refer the reader to the book of Grötschel et al.
+[33].
 Indeed, the practical relevance of $\theta(G)$ is no less.
-On one side, it often turns out to be a significantly stronger upper bound than those from linear relaxations [28, 39].
-On the other side, the structure of the associated SDP relaxation often mitigates numerical difficulties that typically affect SDP algorithms.
-As a consequence, $\theta(G)$ realizes a very good trade-off between the quality of the upper bound and computational effort, which makes it attractive to be incorporated within branch-and-bound algorithms, as studied by, e.g., Wilson [67] and more recently by Gaar et al.
+On one side, it often turns out to be a significantly stronger upper bound than those from linear relaxations, as documented in [28, 39].
+On the other side, the structure of the associated SDP relaxation often mitigates numerical difficulties that typically affects SDP algorithms.
+As a consequence, $\theta(G)$ realizes a very good trade-off between the quality of the upper bound and computational effort, which makes it attractive to incorporate it within branch-and-bound algorithms, as studied by, e.g., Wilson [67] and more recently by Gaar et al.
 [25].
 In a successive landmark paper Lovász and Schrijver [43] introduced the $N_{+}(\cdot)$ lift-and-project operator, which can be applied to the LP relaxation of any 0-1 LP and returns a stronger SDP relaxation.
-Its application to $\operatorname{FRAC}(G)$ has also been intensely investigated and has resulted tighter than the Lovász theta relaxation.
-From a theoretical perspective, its strength has been documented by showing implications of (additional) facet defining inequalities of the stable set polytope [43, 27].
-Moreover, as noted in [7, 8], it provides an effective tool to identify other classes of graphs for which the SSP can be solved in polynomial time.
+Its application to $\operatorname{FRAC}(G)$ has also been intensively investigated and has resulted in a tighter relaxation than the Lovász theta relaxation.
+From a theoretical perspective, its strength has been documented by showing implications of (additional) facet defining inequalities of the stable set polytope (Giandomenico and Letchford [27], Lovász and Schrijver [43]).
+Moreover, Bianchi et al.
+[7, 8] showed that it provides an effective tool to identify other classes of graphs for which the SSP can be solved in polynomial time.
 Recently, graph classes yielding stable set polytopes with high rank with respect to $N_{+}(\cdot)$ have been studied in [1].
 From the computational point of view, things are somewhat more involved.
-Experiments conducted with $N_{+}(\operatorname{FRAC}(G))$ [21, 12] and related variants [23, 34] show that these relaxations may yield stronger upper bounds than $\theta(G)$ but at the price of considerably increasing the computational effort.
-In fact, solving these large-scale SDPs requires specialized methods and large instances can hardly be managed.
+Experiments conducted with $N_{+}(\operatorname{FRAC}(G))$ by Dash [21], Burer and Vandenbussche [12] and with related variants by Dukanovic and Rendl [23], Gruber and Rendl [34], show that these relaxations may yield stronger upper bounds than $\theta(G)$ but at the price of considerably increasing the computational effort.
+In fact, solving these large-scale SDPs requires specialized methods, and large instances can hardly be managed.
 Other methods, based on strengthening SDP relaxations by linear inequalities not related to the $N_{+}(\cdot)$ operator, have also been presented by Gaar and Rendl [24] and Locatelli [41].
-These achieve strong bounds but still require a significant computational effort.
+These achieve strong bounds but still require significant computational effort.
 Overall, we currently face a dichotomic picture, where even minimal progress from the Lovász $\theta$ bound may be paid with a very high computational burden.
@@
 The set of integers $\{1,...,k\}$ is denoted by $[k]$.
-The complement of a graph $G=(V,E)$ is the graph ${\bar{G}}=\big{(}V,{\bar{E}}\big{)}$, where ${\bar{E}}=\big{\{}\{i,j\}\subset V:\,\{i,j\}\in(V\times V)\setminus E\big{\}}$.
-Given a vertex $i\in V$, we let $\Gamma_{G}(i)$ denote the set of neighbours of $i$ in $G$, i.e., $\Gamma_{G}(i)=\big{\{}j\in V\setminus\{i\}:\{i,j\}\in E\big{\}}$.
+The complement of a simple graph $G=(V,E)$ is the graph ${\bar{G}}=\big(V,{\bar{E}}\big)$, where ${\bar{E}}=\big\{\{i,j\}\subset V:\,\{i,j\}\in(V\times V)\setminus E,i\not=j\big\}$.
+Given a vertex $i\in V$, we let $\Gamma_{G}(i)$ denote the set of neighbors of $i$ in $G$, i.e., $\Gamma_{G}(i)=\big\{j\in V\setminus\{i\}:\{i,j\}\in E\big\}$.
 When no confusion arises, we let $\Gamma_{G}(i)=\Gamma(i)$.
-Finally, given a vertex $S\subseteq V$, we let $G[S]=\big{(}S,E(S)\big{)}$ denote the subgraph of $G$ induced by $S$, and $r(S)$ denote its stability number $\alpha\big{(}G[S]\big{)}$ (the so-called rank of $S$).
+Finally, given a set $S\subseteq V$, we let $G[S]=\big(S,E(S)\big)$ denote the subgraph of $G$ induced by $S$, and $r(S)$ denote its stability number $\alpha\big(G[S]\big)$ (the so-called rank of $S$).
 We also let ${\cal S}_{n}$ be the set of real symmetric square matrices of order $n$ and ${\cal S}^{+}_{n}~\subset~{\cal S}_{n}$ denotes the cone of those which are positive semidefinite (PSD).
 PSDness of a matrix $Y\in{\cal S}^{+}_{n}$ is also denoted by $Y\succeq 0$.
-Given any square matrix $X\in\mathbb{R}^{n\times n}$ we denote by $(X)_{ij}$ the element at the $i$-th row and the $j$-th column, by $\operatorname{diag}(X)\in\mathbb{R}^{n}$ its main diagonal and by $\operatorname{rank}(X)$ its rank.
+Given any square matrix $X\in\mathbb{R}^{n\times n}$ we denote by $X_{ij}$ the element at the $i$-th row and the $j$-th column, by $\operatorname{diag}(X)\in\mathbb{R}^{n}$ its main diagonal and by $\operatorname{rank}(X)$ its rank.
 We denote by $\left\langle X,Y\right\rangle=\operatorname{tr}(XY)$ the standard inner product in ${\cal S}_{n}$, where $\operatorname{tr}(A)$ is the trace of the square matrix $A$.
-At last, we denote with $J$ the all-ones square matrix and $e$ the all-ones vector; the dimensions should be clear from the context where not explicit.
+At last, we denote by $J$ the all-ones square matrix and $e$ the all-ones vector; the dimensions should be clear from the context where not explicit.
+Given $a\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$, the linear inequality $a^{\top}x\leq b$ is valid for $\operatorname{STAB}(G)$ if $\left\{x\in\mathbb{R}^{n}\mid a^{\top}x\leq b\right\}\supseteq\operatorname{STAB}(G).$
 2.1 The lift-and-project operator of Lovász and Schrijver
@@
 Let us consider a 0-1 linear program of the form
-$\text{max}\left\{c^{\top}x\ |\ Ax\leq b,\ x_{i}\in\{0,1\}\text{ for }i\in I=\{1,\ldots,n\}\right\},$ (IP)
+$\text{max}\left\{c^{\top}x\ |\ Ax\leq b,\ x_{i}\in\{0,1\}\text{ for }i\in[n]\right\},$ (IP)
 where $c\in\mathbb{R}^{n}$, the inequality system $Ax\leq b$ has $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$ and contains the inequalities $0\leq x_{i}\leq 1$, for $i\in[n]$.
@@
 Generate the quadratic inequalities
-$x_{i}((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m],$ (2)
-$(1-x_{i})((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m].$ (3)
+$x_{i}\cdot((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m],$ (2)
+$(1-x_{i})\cdot((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m].$ (3)
 which are valid for $P$ as $x_{i}\geq 0$ and $1-x_{i}\geq 0$ for all $x\in P$.
-These inequalities can be linearized through the introduction of the matrix variable $X=xx^{\top}$ and rewriting (2) and (3) as:
+These inequalities can be linearized through the introduction of the matrix variable $X=xx^{\top}$ and  (2) and (3) can be rewritten as
 $\sum_{j\in I}a^{k}_{j}X_{ij}-b^{k}x_{i}$ $\leq 0\quad i\in[n],k\in[m],$ (4)
@@
 The projection of $M_{+}(P)$ onto the original $x$-space is
-${N_{+}(P)}=\left\{x\in\mathbb{R}^{n}\ \Big{|}\ \exists\ Y\in M_{+}(\text{P})\text{ with }\begin{pmatrix}1\\ x\end{pmatrix}=\text{diag}(Y)\right\}.$
-[…]
+${N_{+}(P)}=\left\{x\in\mathbb{R}^{n}\ \Big|\ \exists\ Y\in M_{+}(P)\text{ with }\begin{pmatrix}1\\ x\end{pmatrix}=\text{diag}(Y)\right\}.$
+Lovász and Schrijver [43] showed that $N_{+}(P)$ is contained in $P$ and is a valid relaxation for $P_{I}$, i.e.
+$P_{I}\subseteq{N_{+}(P)}\subseteq P$ holds.
+For the ease of presentation, in the reminder of this paper we will refer to both $M_{+}(P)$ and $N_{+}(P)$ as a relaxation of $P_{I}$, despite the former “lives” in the lifted variables space.
+Furthermore, note that all the linear inequalities $Ax\leq b$ are trivially satisfied by points in $M_{+}(P)$.
 Remark 1.
 Let $(a^{k})^{\top}x\leq b^{k}$, for $k\in[m]$, be an inequality in the linear system defining $P$, then from the linear inequalities defining $M_{+}(P)$ one can sum (4) and (5), for any $i\in[n]$, to recover the original constraint.
-Moreover, Bodur et al.
-[9] noted that the definition of $M_{+}(P)$ and $N_{+}(P)$ does not depend on the representation of $P$ but only on the set of points in $P$.
-Hence, the following property holds:
+Moreover, Bodur et al. [9] noted that the definition of $M_{+}(P)$ and $N_{+}(P)$ does not depend on the representation of $P$ but only on the set of points in $P$. Hence, the following property holds
 Property 1.
-Given two polytopes $P,P^{\prime}$ with $P^{\prime}\subseteq P$ then ${M_{+}(P^{\prime})}\subseteq{M_{+}(P)}$
+Given two polytopes $P,P^{\prime}$ with $P^{\prime}\subseteq P$ then ${M_{+}(P^{\prime})}\subseteq{M_{+}(P)}$.
 Additionally, given any polytope $P$ containing $\operatorname{STAB}(G)$, Lovász and Schrijver [43] provided necessary conditions for an inequality $a^{\top}x\leq b$ to be valid for $N_{+}(P)$.
@@
 3.1 Linear relaxations of the SSP
-Several classes of valid inequalities for $\operatorname{STAB}(G)$ have been identified since early seventies [33, 11].
-Among them, clique inequalities $\sum_{i\in C}x_{i}\leq 1$, where $C\subseteq V$ is a maximal clique (set of pairwise adjacent vertices), induce facets of STAB$(G)$ [53].
+Several classes of valid inequalities for $\operatorname{STAB}(G)$ have been identified since early seventies, as is illustrated in [11, 33].
+Among them, clique inequalities $\sum_{i\in C}x_{i}\leq 1$, where $C\subseteq V$ is a maximal clique (set of pairwise adjacent vertices), induce facets of STAB(G)(G) [53].
 As customary in the literature, we define the polytope
 ${\operatorname{QSTAB}(G)}=\{x\in[0,1]^{|V|}:\sum_{i\in C}x_{i}\leq 1,\;C\in{\cal K}\}$
-where ${\cal K}$ denotes the collection of all maximal cliques of $G$.
-As clique inequalities clearly imply edge inequalities, STAB$(G)\subseteq$ $\operatorname{QSTAB}(G)$subseteq FRAC$(G)$ and inclusion is generally strict [33].
-Clique inequalities in $G$ can be exponential in number, and the associated separation problem is NP-hard in the strong sense.
-However, effective separation heuristics make this class of inequalities computationally easy to handle, as shown in [45].
-The 0-1 linear program
-$\max$ $\sum\limits_{i\in V}w_{i}x_{i}$
-s.t. $\sum\limits_{i\in C}x_{i}\leq\,1\quad(C\in{\mathcal{C}})$ (8)
-$x_{i}\in\{0,1\}\quad(i\in V).$
-is a valid formulation for the SSP if ${\mathcal{C}}\subseteq{\cal K}$ is any collection of maximal cliques covering all the edges of $G$ (i.e., both the endpoints of every edge are contained in at least one clique of $\mathcal{C}$).
+[…]
+$\max$ $\sum\limits_{i\in V}x_{i}$
+s.t. $\sum\limits_{i\in C}x_{i}\leq\,1\quad C\in{\mathcal{C}}$ (8)
+$x_{i}\in\{0,1\}\quad i\in V.$
+is a valid formulation for the SSP , for all ${\mathcal{C}}\in{\cal Q}$.
 Extending the standard notation, we denote by $\operatorname{QSTAB}(G,\mathcal{C})$ the polytope associated with its continuous relaxation.
-It’s not hard to see that ${\operatorname{QSTAB}(G)}\subseteq{\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC(G)}$.
-$\operatorname{QSTAB}(G,\mathcal{C})$ has a number of clique inequalities bounded by $|E|$ and it typically contains considerably fewer inequalities than $\operatorname{FRAC(G)}$.
-In fact, formulations based on a greedily computed minimum size collection $\mathcal{C}$ are competitive in practice, as shown in several studies (see, e.g., Letchford et al.
+Unless stated otherwise, the results presented in the remainder of the paper are valid for any ${\mathcal{C}}\in{\cal Q}$.
+It is not hard to see that ${\operatorname{QSTAB}(G)}\subseteq{\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$.
+If ${\mathcal{C}}\in{\cal Q}$ is a minimal edge clique cover, then $|\mathcal{C}|\leq|E|$.
+In practice, formulations based on greedily computed edge clique covers $\mathcal{C}$, which are made up of maximal cliques, contain considerably fewer inequalities than those describing $\operatorname{FRAC(G)}$, and also provide a good approximation of $\operatorname{QSTAB}(G)$}, as shown in several studies (see, e.g., Letchford et al.
 [40] and references therein).
@@
 [28] for a survey).
-A well-known example arises when $S=H$, $|H|\geq 5$, induces a chordless cycle of odd cardinality, yielding the odd-hole inequality $\sum_{i\in H}x_{i}\leq\lfloor\frac{|H|}{2}\rfloor$.
-Among other subgraphs, we recall webs and antiwebs introduced in [63].
+A well-known example arises when $S=H$, $|H|\geq 5$, induces a chordless cycle of odd cardinality, yielding the odd-hole inequality $\sum_{i\in H}x_{i}\leq\lfloor|H|/2\rfloor$.
+Among other subgraphs, we recall webs and antiwebs introduced by Trotter Jr [63].
 Let $p$ and $q$ be integers satisfying $p>2q+1$ and $q>1$, and use arithmetic modulo $p$.
-A $(p,q)$--web is a graph with vertex set $\{1,...,p\}$ and with edges from $i$ to $\{i+q,\ldots,i-q\}$, for every $1\leq i\leq p$.
-A $(p,q)$--antiweb is the complement of a $(p,q)$--web.
-The web inequalities take the form $\sum_{i\in W}x_{i}\leq q$ for every vertex set $W$ inducing a $(p,q)$-web, and the antiweb inequalities take the form $\sum_{i\in AW}x_{i}\lfloor p/q\rfloor$ for every vertex set $AW$ inducing a $(p,q)$-antiweb.
-A different modeling rationale has been pursued by Della Croce and Tadei [22] who considered the surrogate relaxation of formulation (1) obtained by summing up, for every $i\in V$, the edge inequalities over all $j\in\Gamma(i)$.
+A $(p,q)$-web is a graph with vertex set $\{1,...,p\}$ and with edges from $i$ to $\{i+q,\ldots,i-q\}$, for every $1\leq i\leq p$.
+A $(p,q)$-antiweb is the complement of a $(p,q)$-web.
+The web inequalities take the form $\sum_{i\in W}x_{i}\leq q$ for every vertex set $W$ inducing a $(p,q)$-web, and the antiweb inequalities take the form $\sum_{i\in AW}x_{i}\leq\lfloor p/q\rfloor$ for every vertex set $AW$ inducing a $(p,q)$-antiweb.
+A different modeling rationale has been pursued by Della Croce and Tadei [22] who considered the surrogate relaxation of formulation based on (1) obtained by summing up, for every $i\in V$, the edge inequalities over all $j\in\Gamma(i)$.
 This yields the so called nodal inequality $\sum_{j\in\Gamma(i)}x_{j}+|\Gamma(i)|x_{i}\,\leq\,|\Gamma(i)|$.
 Murray and Church [48] observed that this relaxation can be strengthened by replacing $|\Gamma(i)|$ with any value $r_{i}\geq\alpha(G[\Gamma(i)])$, which returns a class of 0-1 linear programs of the form
-$\max$ $\,\,\sum\limits_{i\in V}w_{i}x_{i}$
-s.t. $\sum\limits_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i}\quad(i\in V)$ (9)
-$x_{i}\in\{0,1\}\quad(i\in V).$
+$\max$ $\,\,\sum\limits_{i\in V}x_{i}$
+s.t. $\sum\limits_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i}\quad i\in V$ (9)
+$x_{i}\in\{0,1\}\quad i\in V.$
 The strongest formulation, corresponding to $r_{i}=\alpha(G[\Gamma(i)])$, has been extensively investigated in [40].
@@
 Then, we introduce the following notation
-${\operatorname{NOD}(G,r)}=\left\{x\in[0,1]^{|V|}:\sum_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i},\hskip 5.0pti\in V\right\}$
-to denote the polytope defined from (9) (and non-negativity), where $r\in\{\Gamma,\theta,\alpha\}$ specifies the coefficient used.
-Clearly, we have ${\operatorname{NOD}(G,\alpha)}\subseteq{\operatorname{NOD}(G,\theta)}\subseteq{\operatorname{NOD}(G,\Gamma)}$.
+${\operatorname{NOD}(G,r)}=\left\{x\in[0,1]^{|V|}:\sum_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i},\enskip i\in V\right\}$
+to denote the polytope defined from (9) (and non-negativity), where $r\in\{\Gamma,\theta,\alpha\}$ specifies the coefficient used. Clearly, we have
+${\operatorname{NOD}(G,\alpha)}\subseteq{\operatorname{NOD}(G,\theta)}\subseteq{\operatorname{NOD}(G,\Gamma)}.$
 It is worth noting that nodal inequalities in general do not imply edge inequalities, even in the case $r=\alpha$.
@@
 3.2.1 The Lovász theta relaxation
-The Lovász theta relaxation has been introduced in 1979, yet we refer to the presentation of [43].
+The Lovász theta relaxation [42] has been introduced in 1979, yet we refer to the presentation of Lovász and Schrijver [43].
 Introduce the quadratic variable $X_{ij}$, representing the product $x_{i}x_{j}$ for all $\{i,j\}\in V\times V$ and let $X=xx^{\top}$ be the associated symmetric matrix of order $n=|V|$.
-An upper bound to $\alpha(G)$ is given by
-$\theta(G,w)=\max$ $\;\sum_{i\in V}w_{i}x_{i}$ (SDP-$\theta$)
-s.t. $\;X_{ii}=x_{i},\quad i\in V,$
+An upper bound on $\alpha(G)$ is given by
+$\theta(G)=\max$ $\;\sum_{i\in V}x_{i}$
+s.t. $\;X_{ii}=x_{i},\quad i\in V,$ (SDP-$\theta$)
 $\;X_{ij}=0,\quad\{i,j\}\in E,$
-$\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\succeq 0.$ (10)
-We simply denote it by $\theta(G)$ in the unweighted case.
-The projection $\operatorname{TH}(G)$ of the feasible region of (SDP-$\theta$) onto the subspace of the $x_{i}$ variables is known as the theta body, a convex but not polyhedral set in general.
+$\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\succeq 0.$
+The projection $\operatorname{TH}(G)$ of the feasible region of (SDP-$\theta$) onto the subspace of the $x_{i}$ variables is known as the theta body, a convex but not a polyhedral set in general.
 Remarkably, Grötschel et al.
@@
 This implies that there exists a polynomial-time separation algorithm for a class of inequalities which includes all clique inequalities.
-A thorough comparison between the original formulation in [42] and (SDP-$\theta$) is illustrated in [26].
+A thorough comparison between the original formulation in [42] and (SDP-$\theta$) is provided by Galli and Letchford [26].
 $\theta(G)$ can be computed in polynomial time up to an arbitrary precision and often provides a strong bound to the stability number $\alpha(G)$, typically better than those obtained from linear relaxations see, e.g., [28, 30, 38, 69].
-Another interesting feature of (SDP-$\theta$) concerns with its computational behavior.
+Another interesting feature of (SDP-$\theta$) concerns its computational behavior.
 In fact, (SDP-$\theta$) is handled better than similarly sized unstructured SDPs by general SDP solvers.
@@
 From a theoretical viewpoint, Busygin and Pasechnik [13] showed that, unless P=NP, no polynomially computable upper bound to $\alpha(G)$ which is provably smaller than $\theta(G)$ can be found.
-In this context, several stronger relaxations have been obtained by adding linear inequalities to (SDP-$\theta$), as reviewed in the following subsections.
+However, several stronger relaxations have been obtained by adding linear inequalities to (SDP-$\theta$), as reviewed in the following subsections.
 3.2.2 The Schrijver relaxation
-Schrijver [59] observed that the PSD condition (10) does not imply the non-negativity on $X$. Hence, by adding the inequalities
-$X_{ij}\geq 0,\quad\forall\{i,j\}\in\bar{E},$ (SDP-$\theta^{+}$)
-to (SDP-$\theta$), one yields a model that we denote as (SDP-$\theta^{+}$) and the corresponding upper bound as $\theta^{+}(G)$.
+Schrijver [59] observed that the PSD condition in (SDP-$\theta$) does not imply the non-negativity on $X$. Hence, by adding the inequalities
+$X_{ij}\geq 0,\quad\{i,j\}\in\bar{E}$ (SDP-$\theta^{+}$a)
+to (SDP-$\theta$), one yields a model that we denote as (3.2.2) and the corresponding upper bound as $\theta^{+}(G)$.
 Accordingly, $\operatorname{TH}^{+}(G)$ denotes the convex set obtained by its projection onto the subspace of $x_{i}$ variables.
@@
 The Lovász-Schrijver relaxation for the SSP is obtained by applying the operator $N_{+}(\cdot)$ to $\operatorname{FRAC}(G)$.
-$N_{+}(\operatorname{FRAC}(G))$ corresponds to (SDP-$\theta^{+}$) plus the following linear inequalities:
-$X_{ij}$ $\geq 0,$ $\{i,j\}\in\bar{E},$ (11)
-$X_{ik}+X_{jk}$ $\leq x_{k},$$\{i,j\}\in E,k\neq i,j,$ (12)
-$x_{i}+x_{j}+x_{k}$ $\leq 1+X_{ik}+X_{jk},$$\{i,j\}\in E,k\neq i,j.$ (13)
+$N_{+}(\operatorname{FRAC}(G))$ corresponds to (SDP-$\theta$) plus the following linear inequalities:
+$X_{ij}$ $\geq 0,$ $\{i,j\}\in\bar{E},$ (11a)
+$X_{ik}+X_{jk}$ $\leq x_{k},$ $\{i,j\}\in E,k\neq i,j,$ (11b)
+$x_{i}+x_{j}+x_{k}$ $\leq 1+X_{ik}+X_{jk},$ $\{i,j\}\in E,k\neq i,j.$ (11c)
 Lovász and Schrijver [43] assessed the theoretical strength of $N_{+}(\operatorname{FRAC}(G))$ by showing that it satisfies several well-known classes of valid inequalities for $\operatorname{STAB}(G)$.
@@
 Giandomenico and Letchford [27] showed that, in addition, it satisfies all web inequalities.
-Optimizing $\sum_{i\in V}x_{i}$ over $N_{+}(\operatorname{FRAC}(G))$ returns an upper bound to $\alpha(G)$ that we denote as $\lambda(G)$.
+Optimizing $\sum_{i\in V}x_{i}$ over $N_{+}(\operatorname{FRAC}(G))$ returns an upper bound on $\alpha(G)$ that we denote as $\lambda(G)$.
 From a more general perspective, it is interesting to observe the key role of the positive semidefiniteness constraint.
@@
 The first related computational results, obtained by a lift-and-project cutting plane method, appear in [3].
-Results on optimising directly over $M_{+}(\operatorname{FRAC}(G))$ are illustrated in [21] and, more recently, in [12] using a specialised augmented Lagrangian method.
-The resulting upper bounds to $\alpha(G)$ may be considerably better than $\theta^{+}(G)$ but at the expense of a large increase in running times.
+Results on optimizing directly over $M_{+}(\operatorname{FRAC}(G))$ are illustrated in Dash [21] and, more recently, in Burer and Vandenbussche [12] using a specialized augmented Lagrangian method.
+The resulting upper bounds on $\alpha(G)$ may be considerably better than $\theta^{+}(G)$ but at the expense of a large increase in running times.
 In practice, this relaxation turns out to be hardly tractable for graphs of medium/large size.
 3.2.4 Further SDP relaxations
-Variants of the previous relaxations are presented in [34, 23].
+Variants of the previous relaxations are presented in Dukanovic and Rendl [23] and in Gruber and Rendl [34].
 In the former, extra triangle inequalities are added to $N_{+}(\operatorname{FRAC}(G))$ to obtain a very strong relaxation.
-In the latter, relaxations are specialised with respect to graph density to improve efficiency.
-We refer to [26] for an exhaustive review of these methods.
+In the latter, relaxations are specialized with respect to graph density to improve efficiency.
+We refer to Galli and Letchford [26] for an exhaustive review of these methods.
 Two SDP-related methodologies are also presented in the literature.
-In [24] strong SDP relaxations of several combinatorial optimization problems are investigated by exploiting the exact subgraph inequalities.
+In Gaar and Rendl [24], strong SDP relaxations of several combinatorial optimization problems are investigated by exploiting the exact subgraph inequalities.
 Despite the power of this technique, the experience reveals that strengthening the natural SDP relaxation of the SSP is typically harder than other problems, e.g., max-cut, and confirms that improving the $\theta$ bound towards $\alpha(G)$, even by a small amount, is actually challenging.
-Another effective method (presented in the context of the clique number) based on adding non-valid inequalities has been developed in [41].
+Another effective method (presented in the context of the clique number) based on adding non-valid inequalities has been developed in Locatelli [41].
 This turns out to achieve quite strong bounds for both structured and random graphs with up to 150 vertices.
@@
 Let us first observe that all of them contain variable bound constraints, the lifting step of which consists in multiplying $0\leq x_{i}\leq 1$ by $x_{j}$ and $(1-x_{j})$, for $i,j\in V,i<j$:
-$x_{i}x_{j}\rightarrow$ $\quad X_{ij}\geq 0,$ (14)
-$x_{i}(1-x_{j})\rightarrow$ $\quad x_{i}\geq X_{ij},$ (15)
-$(1-x_{i})x_{j}\rightarrow$ $\quad x_{j}\geq X_{ij},$ (16)
-$(1-x_{i})(1-x_{j})\rightarrow$ $\quad x_{i}+x_{j}-X_{ij}\leq 1.$ (17)
-The linear inequalities (14)–(17) are known as McCormick [46] inequalities and describe the convex hull of the set $\{(x_{i},x_{j},y_{ij})\in\{0,1\}^{3}:y_{ij}=x_{i}x_{j}\}$.
+$x_{i}x_{j}\rightarrow$ $\quad X_{ij}\geq 0,$ (12a)
+$x_{i}(1-x_{j})\rightarrow$ $\quad x_{i}\geq X_{ij},$ (12b)
+$(1-x_{i})x_{j}\rightarrow$ $\quad x_{j}\geq X_{ij},$ (12c)
+$(1-x_{i})(1-x_{j})\rightarrow$ $\quad x_{i}+x_{j}-X_{ij}\leq 1.$ (12d)
+The linear inequalities (12a)–(12d) are known as McCormick [46] inequalities and describe the convex hull of the set $\{(x_{i},x_{j},y_{ij})\in\{0,1\}^{3}:y_{ij}=x_{i}x_{j}\}$.
 It is easy to see that they are not implied by the PSD condition (6) and should be explicitly added to the relaxations illustrated below unless implied by other constraints.
@@
 The lifting step consists in multiplying each clique inequality $\sum_{i\in C}x_{i}\leq 1$, $C\in{\cal K}$, by $x_{j}$ and $(1-x_{j})$, for all $j\in V$.
-Then, adding the condition $X_{ii}=x_{i}$, one obtains:
-$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C\setminus\{j\}}X_{ij}\leq 0,$ if $j\in C$, (18)
-$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C}X_{ij}-x_{j}\leq 0,$ if $j\not\in C$, (19)
-$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}x_{i}-\sum\limits_{i\in C\setminus\{j\}}X_{ij}\leq 1,$ if $j\in C$, (20)
-$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}\leq 1,$ if $j\not\in C$. (21)
-Inequalities (19), called clique-variable inequalities, have been introduced in [28]. One can observe that
+Then, adding the condition $X_{ii}=x_{i}$, one obtains
+$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C\setminus\{j\}}X_{ij}\leq 0,$ if $j\in C$, (13a)
+$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C}X_{ij}-x_{j}\leq 0,$ if $j\not\in C$, (13b)
+$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}x_{i}-\sum\limits_{i\in C\setminus\{j\}}X_{ij}\leq 1,$ if $j\in C$, (13c)
+$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}\leq 1,$ if $j\not\in C$. (13d)
+Inequalities (13b), called clique-variable inequalities, have been introduced in [28]. One can observe that
   1. 1.
-    Constraints (18) imply $X_{ij}=0$ for all $\{i,j\}\in E$ and constraints (20) reduce to the corresponding clique inequality $\sum_{i\in C}x_{i}\leq 1,C\in{\cal K}$, which is already implied from (19) and (21) by Remark 1;
+    Constraints (13a) imply $X_{ij}=0$ for all $\{i,j\}\in E$ and constraints (13c) reduce to the corresponding clique inequality $\sum_{i\in C}x_{i}\leq 1,C\in{\cal K}$, which is already implied from (13b) and (13d) by Remark 1;
   2. 2.
-    $X_{ij}=0$ for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (19) imply inequalities (15) and (16);
+    $X_{ij}=0$ for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (13b) imply inequalities (12b) and (12c);
   3. 3.
-    Constraints (18), (19) and (21) imply (17).
+    Constraints (13a), (13b) and (13d) imply (12d).
 Then, the description of $M_{+}(\operatorname{QSTAB}(G))$ reduces to
-$\begin{split}{M_{+}(\operatorname{QSTAB}(G))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ \sum\limits_{i\in C}X_{ij}-x_{j}&\leq 0,\quad C\in\mathcal{K},j\in V\setminus C,\\ \sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}&\leq 1,\quad C\in\mathcal{K},j\in V\setminus C,\\ x_{i}&=X_{ii},\quad i\in V,\\ X_{ij}&=0,\quad\{i,j\}\in E,\\ X_{ij}&\geq 0,\quad\{i,j\}\in{\bar{E}}\\ x_{i}&\geq 0,\quad i\in V\bigg{\}}.\end{split}$ (22)
-Moreover, one has:
+$\begin{split}{M_{+}(\operatorname{QSTAB}(G))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ \sum\limits_{i\in C}X_{ij}-x_{j}&\leq 0,\quad C\in\mathcal{K},j\in V\setminus C,\\ \sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}&\leq 1,\quad C\in\mathcal{K},j\in V\setminus C,\\ x_{i}&=X_{ii},\quad i\in V,\\ X_{ij}&=0,\quad\{i,j\}\in E,\\ X_{ij}&\geq 0,\quad\{i,j\}\in{\bar{E}}\\ x_{i}&\geq 0,\quad i\in V\bigg\}.\end{split}$
+Moreover, one has
 Theorem 1.
@@
 Proof Since ${\operatorname{QSTAB}(G)}\subseteq\operatorname{FRAC}(G)$, the statement follows from Property 1. ∎
-We denote by $\mu(G)=\max\left\{\sum_{i\in V}x_{i}:\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\in{M_{+}(\operatorname{QSTAB}(G))}\right\}$ the resulting upper bound to $\alpha(G)$.
-The following example shows that the containment ${M_{+}(\operatorname{QSTAB}(G))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$ can be strict.
+We denote the resulting upper bound on $\alpha(G)$ by
+$\mu(G)=\max\left\{\sum_{i\in V}x_{i}:\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\in{M_{+}(\operatorname{QSTAB}(G))}\right\}$
+The following example shows that the containment of Theorem 1 can be strict.
 Example 1.
@@
 Therefore, $x_{F}\in{N_{+}(\operatorname{FRAC}(G))}$ lies outside $N_{+}(\operatorname{QSTAB}(G))$.
-The example indeed suggests a general result. In fact, we can show that, unlike $N_{+}(\operatorname{FRAC}(G))$, relaxation $N_{+}(\operatorname{QSTAB}(G))$ implies the class of antiweb inequalities.
+The example indeed suggests a general result.
+In fact, we can show that, unlike $N_{+}(\operatorname{FRAC}(G))$, the relaxation $N_{+}(\operatorname{QSTAB}(G))$ implies the class of antiweb inequalities.
 Theorem 2.
-$N_{+}(\operatorname{QSTAB}(G))$ implies antiweb inequalities
+$N_{+}(\operatorname{QSTAB}(G))$ implies antiweb inequalities.
 Proof Let $G$ be the $(p,q)$-antiweb graph with $V=\left\{1,\ldots,p\right\}$.
@@
 Note that, by definition, any consecutive $q$ nodes in $G$ form a clique.
-By denoting $\alpha=\left\lfloor\frac{p}{q}\right\rfloor$, the $(p,q)$-antiweb inequality is
+By denoting $\alpha=\left\lfloor p/q\right\rfloor$, the $(p,q)$-antiweb inequality is
 $\sum\limits_{i\in V}x_{i}\leq\alpha.$
-Given any $i\in V$, we denote with $G\setminus i$ the subgraph induced by the vertex set
+Given any $i\in V$, we denote by $G\setminus i$ the subgraph induced by the vertex set
 $V(G\setminus i)=V\setminus(\left\{i\right\}\cup\Gamma(i))=\left\{i+q,\ i+q+1,\dots,\ i-q-1,\ i-q\right\}.$
@@
 is valid for $\operatorname{QSTAB}(G\setminus i)$}, for all $i\in V$, then the theorem’s statement will follow from Lemma 1.
-Let $i$ be any node in $V$ and consider the following $\alpha-1$ cliques in $G\setminus i$:
-$C^{1}$ =$=\left\{i+q,\ i+q+1,\dots,\ i+2q-1\right\},$
-$C^{2}$ =$=\left\{i+2q,\ i+2q+1,\dots,\ i+3q-1\right\},$
+Let $i$ be any node in $V$ and consider the following $\alpha-1$ cliques in $G\setminus i$
+$C^{1}$ $=\left\{i+q,\ i+q+1,\dots,\ i+2q-1\right\},$
+$C^{2}$ $=\left\{i+2q,\ i+2q+1,\dots,\ i+3q-1\right\},$
 $\vdots$
-$C^{k}$ =$=\left\{i+kq,\ i+kq+1,\dots,i+(k+1)q-1\right\},$
+$C^{k}$ $=\left\{i+kq,\ i+kq+1,\dots,i+(k+1)q-1\right\},$
 $\vdots$
-$C^{\alpha-1}$ =$=\left\{i+(\alpha-1)q,\ i+(\alpha-1)q+1,\dots,\ i-q\right\}.$
-Note that the set $\left\{C^{1},\dots,C^{\alpha-1}\right\}$ forms a partition of $V(G\setminus i)$, i.e., $\bigcup_{k=1,\dots,\alpha-1}C^{k}=V(G\setminus i)$ and $C^{\ell}\cap C^{k}=\emptyset$, for all $\ell,k=1,\dots,\alpha-1$, with $\ell\not=k$.
-Hence, the sum of clique inequalities $\sum_{j\in C^{k}}x_{j}\leq 1$, for $k=1,\dots,\alpha-1$ yield the inequality
+$C^{\alpha-1}$ $=\left\{i+(\alpha-1)q,\ i+(\alpha-1)q+1,\dots,\ i-q\right\}.$
+Note that the set $\left\{C^{1},\dots,C^{\alpha-1}\right\}$ forms a partition of $V(G\setminus i)$, i.e.,
+$\bigcup_{k=1,\dots,\alpha-1}C^{k}=V(G\setminus i)$
+and $C^{\ell}\cap C^{k}=\emptyset$, for all $\ell,k=1,\dots,\alpha-1$, with $\ell\not=k$.
+Hence, the sum of clique inequalities $\sum_{j\in C^{k}}x_{j}\leq 1$, for $k=1,\dots,\alpha-1$ yields the inequality
 $\sum\limits_{i+q\leq j\leq i-q}x_{j}=\sum\limits_{j\in V(G\setminus i)}x_{j}\leq\alpha-1,$
 which is valid for $\operatorname{QSTAB}(G\setminus i)$. By symmetry, this argument is independent of the choice of $i$. ∎
-Theorem 1 completes the theoretical picture for $N_{+}(\operatorname{QSTAB}(G))$ along the line drawn in [43].
+Theorem 1 completes the theoretical picture for $N_{+}(\operatorname{QSTAB}(G))$ along the line drawn by Lovász and Schrijver [43].
 In fact, all of the traditional combinatorial inequalities, namely, clique, odd cycle, odd antihole, odd wheel, web and antiweb inequalities turn out to be satisfied by $N_{+}(\operatorname{QSTAB}(G))$.
@@
 The comparison between $N(K,K)$ and $N_{+}(\cdot)$ was indeed suggested by Lovász and Schrijver [43] as an interesting open issue.
-Another insight of the progress of $N_{+}(\operatorname{QSTAB}(G))$ towards $N_{+}(\operatorname{FRAC}(G))$ stems from the analysis of graph $\hat{G}$ of Figure 1 which has been shown in [6] that $N_{+}(\operatorname{FRAC}(\hat{G}))$ gives rise to a non-polyhedral relaxation.
+Another insight of the progress of $N_{+}(\operatorname{QSTAB}(G))$ towards $N_{+}(\operatorname{FRAC}(G))$ stems from the analysis of graph $\hat{G}$ of Figure 1 which has been shown in [6] that $)N_{+}(\operatorname{FRAC}(\hat{G}$)) gives rise to a non-polyhedral relaxation.
 Here we have $\lambda(\hat{G})=2.146$.
-Now, optimizing over $N_{+}(\operatorname{QSTAB}(\hat{G}))$ one gets $\mu(\hat{G})=2$.
-This implies that $N_{+}(\operatorname{QSTAB}(\hat{G}))$ is actually polyhedral.
+Now, optimizing over $)N_{+}(\operatorname{QSTAB}(\hat{G}$)) one gets the stability number $\mu(\hat{G})=\alpha(\hat{G})=2$.
+This may suggest that $)N_{+}(\operatorname{QSTAB}(\hat{G}$)) is actually polyhedral.
 Figure 1: The $\hat{G}$ graph.
-From a computational perspective, a major issue of $N_{+}(\operatorname{QSTAB}(G))$ is the linear description of $\operatorname{QSTAB}(G)$ that, in general, contains exponentially many inequalities.
-However, as ${\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$, we still have ${M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$ and one can consider the convex set:
+From a computational perspective, a major issue of $N_{+}(\operatorname{QSTAB}(G))$ is that the linear description of $\operatorname{QSTAB}(G)$ , in general, contains exponentially many inequalities.
+To get rid of this issue, we consider $\operatorname{QSTAB}(G,\mathcal{C})$}, for any edge clique cover $\cal C\in\cal Q$.
+The application of the lifting operator to this relaxation yields the convex set
 […]
-Note that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ still implies McCormick inequalities (14)–(17).
-Remarkably, it can be built in polynomial time as a suitable set $\cal C$ can be determined by a greedy algorithm (e.g., the one proposed in Letchford et al.
-[40]), and the number of inequalities describing $\operatorname{QSTAB}(G,\mathcal{C})$ is bounded by $|E|$ (for increasing graph density, it is also considerably smaller than $|E|$).
+Since ${\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$, we have ${M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$.
+Also note that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ still implies McCormick inequalities (12a)–(12d).
+Remarkably, it can be built in polynomial time as a minimal edge clique cover ${\cal C}\in{\cal Q}$ can be determined by a greedy algorithm (e.g., the one proposed in Letchford et al.
+[40]), and the number of inequalities describing $\operatorname{QSTAB}(G,\mathcal{C})$ is at most $|E|$ (for increasing graph density, it is also considerably smaller than $|E|$).
 Optimizing $\sum_{i\in V}x_{i}$ over $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ returns an upper bound that we denote as $\mu(G,{\cal C})$.
@@
 Let us now consider a generic nodal inequality, with $r_{i}\geq\alpha(G[\Gamma(i)])$, for all $i\in V$. The lifting step along with $x_{i}=X_{ii}$ returns:
-$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}X_{ih}\leq 0,$ if $j=i$ (24)
-$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $(1-r_{i})x_{j}+\sum_{h\in\Gamma(i)\setminus j}X_{jh}+r_{i}X_{ij}\leq 0,$ if $j\in\Gamma(i)$ (25)
-$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}X_{jh}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,$ if $j\not\in\Gamma(i)$ (26)
-$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}(x_{h}-X_{ih})+r_{i}x_{i}\leq r_{i},$ if $j=i$ (27)
-$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)\setminus j}(x_{h}-X_{jh})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\in\Gamma(i)$ (28)
-$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}(x_{h}-X_{jh})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\not\in\Gamma(i)$ (29)
+$x_{j}\cdot\left(\sum_{k\in\Gamma(i)}x_{k}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$
+$\rightarrow$ $\sum_{k\in\Gamma(i)}X_{ik}\leq 0,$ if $j=i$ (15a)
+$\rightarrow$ $(1-r_{i})x_{j}+\sum_{k\in\Gamma(i)\setminus j}X_{jk}+r_{i}X_{ij}\leq 0,$ if $j\in\Gamma(i)$ (15b)
+$\rightarrow$ $\sum_{k\in\Gamma(i)}X_{jk}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,$ if $j\not\in\Gamma(i)$ (15c)
+$(1-x_{j})\cdot\left(\sum_{k\in\Gamma(i)}x_{k}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$
+$\rightarrow$ $\sum_{k\in\Gamma(i)}(x_{k}-X_{ik})+r_{i}x_{i}\leq r_{i},$ if $j=i$ (16a)
+$\rightarrow$ $\sum_{k\in\Gamma(i)\setminus j}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\in\Gamma(i)$ (16b)
+$\rightarrow$ $\sum_{k\in\Gamma(i)}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\not\in\Gamma(i)$ (16c)
 One can observe that:
   1. 1.
-    Constraints (24) imply $X_{ij}=0$ for all $\{i,j\}\in E$; this is a relevant fact if one recalls that edge inequalities are not implied by nodal inequalities (9);
+    Constraints (15a) imply $X_{ij}=0$ for all $\{i,j\}\in E$; this is a relevant fact if one recalls that edge inequalities are not implied by nodal inequalities (9);
   2. 2.
-    $X_{ij}=0$, for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (26) imply inequalities (15) and (16);
+    $X_{ij}=0$, for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (15c) imply inequalities (12b) and (12c);
   3. 3.
-    $X_{ij}=0$, for all $\{i,j\}\in E$, inequalities (15) and (16) and constraints (28) imply edge inequalities (and (17) for $\{i,j\}\in E$);
+    $X_{ij}=0$, for all $\{i,j\}\in E$, inequalities (12b) and (12c) and constraints (16b) imply edge inequalities (and (12d) for $\{i,j\}\in E$);
   4. 4.
-    By $X_{ij}=0$, for all $\{i,j\}\in E$, (27) boils down to the original nodal inequality (9) and can be dropped, since Remark 1 can be applied to both pairs (25), (28) and/or (26), (29);
+    By $X_{ij}=0$, for all $\{i,j\}\in E$, (16a) boils down to the original nodal inequality (9) and can be dropped, since Remark 1 can be applied to both pairs (15b), (16b) and (15c), (16c);
   5. 5.
-    $X_{ij}\geq 0$, for all $\{i,j\}\in\bar{E}$, inequalities (15) and (16) and constraints (29) imply (17) for $\{i,j\}\in\bar{E}$.
+    $X_{ij}\geq 0$, for all $\{i,j\}\in\bar{E}$, inequalities (12b) and (12c) and constraints (16c) imply (12d) for $\{i,j\}\in\bar{E}$.
 Thanks to these properties $M_{+}(\operatorname{NOD}(G,r))$ boils down to
-$\begin{split}{M_{+}(\operatorname{NOD}(G,r))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ (1-r_{i})x_{j}+\sum_{h\in\Gamma(i)\setminus j}X_{jh}\leq 0,&\quad i\in V,j\in\Gamma(i),\\ \sum_{h\in\Gamma(i)}X_{jh}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,&\quad i\in V,j\not\in\Gamma(i),\\ \sum_{h\in\Gamma(i)\setminus j}(x_{h}-X_{hj})+r_{i}x_{i}+r_{i}x_{j}\leq r_{i},&\quad i\in V,j\in\Gamma(i),\\ \sum_{h\in\Gamma(i)}(x_{h}-X_{hj})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},&\quad i\in V,j\not\in\Gamma(i),\\ x_{i}=X_{ii},&\quad i\in V,\\ X_{ij}=0,&\quad\{i,j\}\in E,\\ X_{ij}\geq 0,&\quad\{i,j\}\in{\bar{E}},\\ x_{i}\geq 0,&\quad i\in V\bigg{\}}.\end{split}$ (30)
-Then, $M_{+}(\operatorname{NOD}(G,r))$ can be obtained by adding linear inequalities (25), (26), (28), (29) and $X_{ij}\geq 0$, for $\{i,j\}\in\bar{E}$, to (SDP-$\theta$) (it is straightforward to observe that inequalities (25) are implied by Mc-Cormick inequalities for $r_{i}=\Gamma$).
-Therefore, its projection $N_{+}(\operatorname{NOD}(G,r))$ onto the $x$-space is at least as strong as the Schrijver relaxation:
+$\begin{split}{M_{+}(\operatorname{NOD}(G,r))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ (1-r_{i})x_{j}+\sum_{k\in\Gamma(i)\setminus j}X_{jk}\leq 0,&\quad i\in V,j\in\Gamma(i),\\ \sum_{k\in\Gamma(i)}X_{jk}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,&\quad i\in V,j\not\in\Gamma(i),\\ \sum_{k\in\Gamma(i)\setminus j}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}\leq r_{i},&\quad i\in V,j\in\Gamma(i),\\ \sum_{k\in\Gamma(i)}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},&\quad i\in V,j\not\in\Gamma(i),\\ x_{i}=X_{ii},&\quad i\in V,\\ X_{ij}=0,&\quad\{i,j\}\in E,\\ X_{ij}\geq 0,&\quad\{i,j\}\in{\bar{E}},\\ x_{i}\geq 0,&\quad i\in V\bigg\}.\end{split}$
+Then, $M_{+}(\operatorname{NOD}(G,r))$ can be obtained by adding linear inequalities (15b), (15c), (16b), (16c) and $X_{ij}\geq 0$, for $\{i,j\}\in\bar{E}$, to (3.2.2) (it is straightforward to observe that inequalities (15b) are implied by Mc-Cormick inequalities for $r=\Gamma$).
+Therefore, its projection $N_{+}(\operatorname{NOD}(G,r))$ onto the $x$-space is at least as strong as the Schrijver relaxation.
 Theorem 3.
 ${N_{+}(\operatorname{NOD}(G,r))}\subseteq\operatorname{TH}^{+}(G)$.
-Which also implies:
+This result also implies
 Corollary 1.
@@
 Consider the odd-hole with seven vertices of Figure (2(b)).
-According to [43], $N_{+}(\operatorname{FRAC}(G))$ satisfies the facet-defining inequality $\sum_{i=1}^{8}x_{i}\leq 3$ which implies $\lambda(G)=3$.
+According to [43], $N_{+}(\operatorname{FRAC}(G))$ satisfies the facet-defining inequality $\sum_{i\in V}x_{i}\leq 3$ which implies $\lambda(G)=3$.
 Differently, the optimal value from $N_{+}(\operatorname{NOD}(G,r))$ is $\nu(G,\theta)=\nu(G,\alpha)=3.317$ corresponding to solution $(0.47395,\ldots,0.47395)$, which therefore falls outside $N_{+}(\operatorname{FRAC}(G))$.
@@
 Here, $\nu(G,\theta)=\nu(G,\alpha)=\alpha(G)=3$.
-On the contrary, optimizing over $N_{+}(\operatorname{FRAC}(G))$ we obtain solution
+On the contrary, optimizing over $N_{+}(\operatorname{FRAC}(G))$ we obtain the solution
 $x_{F}=(0.31055,\dots,0,31055)$
 with value $3.105$ which is then outside $N_{+}(\operatorname{NOD}(G,r))$, for $r\in\{\theta,\alpha\}$.
-11112222333344445555666677778888999910101010
-(a) $(10,3)$–antiweb
-1111222233334444555566667777
+1122334455667788991010
+(a) (10,3)(10,3)–antiweb
+11223344556677
 (b) 7-hole
@@
 This example also shows that the odd-hole inequalities are not implied by $N_{+}(\operatorname{NOD}(G,r))$ even with $r=\alpha$.
-This appears as a weakness indicator of $N_{+}(\operatorname{NOD}(G,r))$ if one recalls that odd-hole inequalities are contained in the first Chvàtal closure of the edge formulation 1 [53].
-The overall theoretical picture is summarized in Figure 3, and the above results seem to suggest that $N_{+}(\operatorname{QSTAB}(G))$ (as well as $N_{+}(\operatorname{FRAC}(G))$) should be stronger than $N_{+}(\operatorname{NOD}(G,r))$.
+This appears as a weakness indicator of $N_{+}(\operatorname{NOD}(G,r))$ if one recalls that odd-hole inequalities are contained in the first Chvàtal closure of the formulation based on edge inequalities (1) [53].
+The overall theoretical picture is summarized in Figure 3 .
+The results related to the implication of odd-hole and antiweb inequalities seem to suggest that $N_{+}(\operatorname{QSTAB}(G))$ (as well as $N_{+}(\operatorname{FRAC}(G))$) should be stronger than $N_{+}(\operatorname{NOD}(G,r))$.
 Interestingly, the computational analysis will reveal significant supremacy of the latter as graphs get dense, for $r\in\{\theta,\alpha\}$.
-Symbol Relaxation
+Optimal value Relaxation
 $\lambda(G)$ $N_{+}(\operatorname{FRAC}(G))$
@@
 Selection of the SDP solver
-[…]
+Under mild assumptions, semidefinite programs can be solved in polynomial time up to any arbitrary fixed precision [51].
+Intuitively, the computational burden mainly depends on the order $n$ of the matrix variable and the number $m$ of linear constraints.
+However, the density of the latter should also be taken into account as most SDP solvers rely on sparse matrices implementations.
+Interior-point methods (IPMs) [47, 60, 64] provide a good-precision solution in reasonable time for “small” and “medium” SDPs (both in term of $n$ and $m$).
+However, they require to store and factorize large Hessian matrices at each iteration, which may become prohibitive for large SDPs.
+In fact, we preliminarily tested the IPM implemented in the commercial state-of-the-art solver MOSEK [47] and observed that problems with $m\approx 10^{5}$ and $n\approx 200$ are intractable.
+Although this experiment already provided interesting insights on the strength of the relaxations, testing larger graphs is crucial for a reliable experimental picture.
+Therefore, we resorted to Alternating Direction Methods of Multipliers (ADMMs), which, as a variant of Augmented Lagrangian methods, is a popular first-order alternative to IPMs and can scale on much larger SDPs at the price of possibly slowing down the convergence to high-precision solutions.
+For a detailed overview, we refer the reader to [4, 5, 17, 54, 61, 65, 66].
+Furthermore, ADMMs allow the efficient handling of bound constraints (e.g., nonnegativity) on the matrix variable via iterative projections on the cone of positive semidefinite matrices.
+In our case, it enables the computation of $\theta^{+}(G)$ with no substantial additional effort with respect to $\theta(G)$.
+We selected SDPNAL+ proposed in [61], which represents a state-of-the-art ADMM general-purpose SDP solver implemented in MATLAB equipped with efficient subroutines in C included via .mex files, and has proven to be numerically stable in our experiments.
 Model building
-[…]
+The $M_{+}(\cdot)$ operator has been applied to four different compact LP relaxations of the SSP: $\operatorname{QSTAB}(G,\mathcal{C})$ and $\operatorname{NOD}(G,r)$ for $r\in\{\Gamma,\theta,\alpha\}$.
+The construction of these LP models entails different complexity levels.
+$\operatorname{QSTAB}(G,\mathcal{C})$ requires the computation of an edge clique cover $\mathcal{C}\in\mathcal{Q}$ .
+This task has been accomplished by the greedy algorithm presented in Letchford et al.
+[40] and is reported in Algorithm 1.
+Starting from the empty collection $\mathcal{C}=\emptyset$, the heuristic iteratively adds a maximal clique of $G$ to $\mathcal{C}$, until every edge is covered.
+At each iteration, the selected maximal clique $C_{k}$ covers at least one new edge not yet covered by cliques in $\mathcal{C}$ (line 8).
+Consequently, even if it can be shown that the resulting edge clique cover $\mathcal{C}$ may not be minimal, its size is at most $|E|$ and, in practice, comprises significantly fewer cliques [40].
+Algorithm 1 Greedy edge clique cover
+1:Input: Graph $G=(V,E)$
+2:Output: An edge clique cover $\mathcal{C}$ of $G$
+3:Initialize $E^{\prime}\leftarrow E,$ $G^{\prime}\leftarrow(V,E^{\prime})$ $\mathcal{C}\leftarrow\emptyset,\ k\leftarrow 0$
+4:while $E^{\prime}\not$$\emptyset$ do
+5:  $k\leftarrow k+1$
+6:  $u\leftarrow\arg\max_{i\in V}\left\{|\Gamma_{G^{\prime}}(i)|\right\}$
+7:  $v\leftarrow\arg\max_{i\in\Gamma_{G^{\prime}}(u)}\left\{|\Gamma_{G^{\prime}}(i)|\right\}$ $\triangleright$ $E^{\prime}\not=\emptyset\Rightarrow\Gamma_{G^{\prime}}(u)$ is non-empty
+8:  Find a maximal clique $C_{k}$ of $G$ such that $u,v\in C_{k}$
+9:  $\mathcal{C}\leftarrow\mathcal{C}\cup\left\{C_{k}\right\}$
+10:  $E^{\prime}\leftarrow E^{\prime}\setminus E(C_{k})$
+11:end while
+Building $\operatorname{NOD}(G,\Gamma)$ is straightforward, as $r_{i}=|\Gamma(i)|$.
+On the contrary, $\operatorname{NOD}(G,\theta)$ requires the evaluation of $r_{i}=\theta(G[\Gamma(i)])$, for all $i\in V$.
+This corresponds to solve $|V|$ SDPs, namely, those corresponding to (SDP-$\theta$) on the subgraph induced by $\Gamma(i)$.
+In practice, the SDPs are completely decomposed and can be solved in parallel.
+Moreover, $r_{i}$ can be tightened to $\left\lfloor\theta(G[\Gamma(i)])\right\rfloor$ and one can stop the computation as long as the SDP solver returns a valid upper (dual) bound on $\theta(G[\Gamma(i)])$.
+Several approaches to compute a valid dual bound throughout solver’s iterations and a posteriori have been proposed for ADMM algorithms (see, e.g., [17, 36, 66]).
+Their specific application in SDP relaxations of the SSP have been implemented in [4, 5].
+Then, for each iteration of the ADMM implemented in ADMMsolver.py, a valid dual bound is computed and this, in turn, allows stopping the computation of $\theta$ as soon as a “moderate” precision is achieved.
+In Table 2, we report the average/max/min CPU times to calculate coefficients $\theta(G[\Gamma(i)])$ on a subset of the instances used in our computational experiments, whose detailed introduction is deferred to Section 6.1.
+Finally, building $\operatorname{NOD}(G,\alpha)$ is the most challenging case as it requires calculating $\alpha(G[\Gamma(i)])$, for all $i\in V$.
+Despite the theoretical hardness of determining $\alpha(G[\Gamma(i)])$Letchford et al.
+[40] showed that in practice, the computational burden to build $\operatorname{NOD}(G,\alpha)$ is, in fact, accessible for large classes of graphs.
+We refer the interested reader to their discussion; here we mention that even in this case the evaluation of coefficients $r_{i}=\alpha(G[\Gamma(i)])$ can be parallelized and, as in the previous case, we report in Table 2 the average/max/min CPU times to evaluate $\alpha(G[\Gamma(i)])$ via the max-clique solver cliquer [52].
 Graph
@@
 Average 3.93 3.90 3.93 3.93 2.40 3.53 6.14 1.86 7.66 15.95 3.84 15.06 31.25 6.98 1.78 1.09
-$\alpha_{i}$ Min 0.89 1.08 0.68 0.63 $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ 0.01 $<0.01$
-Max 28.87 34.41 26.17 25.91 $<0.01$ $<0.01$ 0.02 0.02 0.02 1.83 2.74 0.07 62.41 108.48 0.04 0.13
-Average 6.00 6.17 5.93 6.04 $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ 0.10 0.10 0.02 1.64 2.78 0.02 0.03
+$\alpha_{i}$ Min 0.89 1.08 0.68 0.63 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 0.01 <0.01<0.01
+Max 28.87 34.41 26.17 25.91 <0.01<0.01 <0.01<0.01 0.02 0.02 0.02 1.83 2.74 0.07 62.41 108.48 0.04 0.13
+Average 6.00 6.17 5.93 6.04 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 <0.01<0.01 0.10 0.10 0.02 1.64 2.78 0.02 0.03
 Table 2: CPU time statistics for computing coefficients $\theta$ and $\alpha$ for (9) on DIMACS graphs.
@@
 Initially, it parses the polytope $P$, provided as input through a standard .lp format text file.
-Then, it construct the symmetric matrices corresponding to constraints (4), (5) and (7).
+Then, it constructs the symmetric matrices corresponding to constraints (4), (5) and (7).
 Finally, the model is exported in .mat file format, ready to be loaded into MATLAB and solved using SDPNAL+.
@@
 and $m\in\mathbb{N}$, $C,A_{1},\ldots,A_{m}\in{\cal S}_{n+1}$, $b\in\mathbb{R}^{m}$.
-In a preliminary experiment we observed that SDPNAL+ converges quickly on (SDP-$\theta^{+}$) while it requires very large (and practically unacceptable) convergence times on $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ relaxations.
+In a preliminary experiment we observed that SDPNAL+ converges quickly on (3.2.2) while it requires very large (and practically unacceptable) convergence times on $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ relaxations.
 This is due to the number of linear constraints in the former, and their density in the latter.
-However, we proved that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ are contained in (SDP-$\theta^{+}$), thus one can resort to a classical cutting plane method to optimize over them.
-The cutting plane algorithm starts optimizing over (SDP-$\theta^{+}$).
+However, we proved that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ are contained in (3.2.2), thus one can resort to a classical cutting plane method to optimize over them.
+The cutting plane algorithm starts optimizing over (3.2.2).
 At the $k$-th iteration, given the optimal solution $Y^{k}\in{\cal S}^{+}_{n+1}$, one can compute
 $V_{i}(Y^{k})=\langle A_{i},Y^{k}\rangle-b_{i},\ i\in[m].$
-[…]
+If $V_{i}(Y^{k})>\varepsilon$, for a “small” $\varepsilon>0$, then the $i$-th constraint is violated by the current solution.
+Let $c\leq m$ be a positive integer parameter, and let $I_{c}(Y^{k})=\left\{i_{1},i_{2},\ldots,i_{c}\right\}$ denote the set of indexes corresponding to the $c$ most violated constraints.
+Then, the constraints $\left\{Y\in{\cal S}_{n+1}\ :\ \langle A_{i},Y\rangle\leq b_{i},\ i\in I_{c}(Y^{k})\right\}$ are added to the current SDP relaxation and a new optimal solution $Y^{k+1}$ is computed.
+The algorithm terminates either when no violated constraints are found or when a tailing off condition is met.
+The detailed algorithm is reported in Algorithm 2.
 Algorithm 2 Cutting-plane scheme
@@
   • (i)
-    Can relaxations based on operator $M_{+}(\cdot)$ provide upper bounds for $\alpha(G)$ that are significantly stronger than $\theta(G)$?
+    Can relaxations based on operator $M_{+}(\cdot)$ provide upper bounds on $\alpha(G)$ that are significantly stronger than $\theta(G)$?
   • (ii)
@@
 As a byproduct, we document the quality of the bound provided by $M_{+}(\operatorname{NOD}(G,\theta))$, that is, the strongest bound which can be computed in polynomial time in our hierarchy.
+Finally, we investigate the extent to which the bound $\mu(G,\mathcal{C})$ is sensitive to the choice of $\mathcal{C}$ and, in particular, the degree to which it can be improved with a more refined selection of the edge clique cover.
 Experiments are run on a computer with Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz, 256GB RAM, under OS Ubuntu 16.04.7 LTS (GNU/Linux 4.15.0-128-generic x86_64).
-After preliminary tuning experiments, the parameters of Algorithm 1 have been set to $\epsilon=10^{-3},\delta=10^{-1}$ and $c=1000$ with 7200 seconds time limit (excluding formulation building time).
+After preliminary tuning experiments, the parameters of Algorithm 2 have been set to $\varepsilon=10^{-3},\delta=10^{-1}$ and $c=1000$ with 7200 seconds time limit (excluding formulation building time).
 The precision of the SDP solver SDPNAL+ is set to $10^{-6}$.
@@
 A Erdös–Rényi graph $G(n,p)$ has $n$ vertices and each edge appears with probability $p\in[0,1]$. We have considered the collection of graphs from [40], generated by
-$n\in\{200,250,300\},\quad p\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}.$
+$n\in\{200,250,300\},\quad p\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\},$
 where, for each combination of $n$ and $p$, five instances with different random seeds have been created for a total of 135 graphs.
 DIMACS graphs
-Graphs from the Second DIMACS Implementation Challenge [37].
-These form the standard benchmark for max-clique algorithms.
+Graphs from the Second DIMACS Implementation Challenge [37] form the standard benchmark for max-clique algorithms.
 We complemented the graphs to convert the max-clique instances into SSP instances (DSJC* graphs are not complemented, as these belong to the “coloring” benchmark set).
 We excluded the c-fat, johnson, and san graphs because the integrality gap is completely closed by linear relaxations, so they do not provide information to our analysis.
-We also dropped keller5 and the largest and p-hat instances (with 1000 and 1500 vertices) as the SDP relaxations turned out to be too large to be handled.
+We also dropped keller5 and the largest p-hat instances (with 1000 and 1500 vertices) as the SDP relaxations turned out to be too large to be handled.
 The final collection includes 38 graphs.
@@
 Table 3 collects gap statistics for Erdös–Rényi graphs, where the first column reports the integrality gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$.
-Tables 4 and 5 report cutting plane iterations, CPU times, and, respectively, the number of added cuts for each class of inequalities.
+Tables 4 and 5 report cutting plane iterations and CPU times, and the number of added cuts for each class of inequalities, respectively.
 Let us first notice that $\lambda(G)\simeq\theta^{+}(G)$ in all cases but $n=200$, $p\in\{0.1,0.9\}$, evidencing a rather disappointing behavior of $M_{+}(\operatorname{FRAC}(G))$.
@@
 $\lambda(G)$ is $6.7\%$.
-This is achieved by a few cutting plane iterations, adding about one to three thousand inequalities of class (19), resulting in an average $80\%$ CPU time increase.
-Interestingly, no inequalities in the class (21) have been detected as violated.
+This is achieved by a few cutting plane iterations, adding about one to three thousand inequalities of class (13b), resulting in an average $80\%$ CPU time increase.
+Interestingly, no inequalities in the class (13d) have been detected as violated.
 This experiment clearly classifies relaxations strength as a function of edge probability $p$ and graph size $n$.
@@
 The outcome is completely reversed when $p\geq 0.5$, where relaxations based on nodal inequalities are more competitive and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$ is the clear winner.
-$\nu(G,\alpha)$ closes $45.3\%$ of the gap left by $\lambda(G)$ and $40.0\%$ of the one left by $\mu(G,{C})$, on average.
+$\nu(G,\alpha)$ closes $45.3\%$ of the gap left by $\lambda(G)$ and $40.0\%$ of the one left by $\mu(G,{\cal C})$, on average.
 To our knowledge, such an improvement on $\theta^{+}(G)$ has rarely been documented in the literature on SDP approaches, even for smaller graphs.
 Another nice evidence is that this advantage of $M_{+}(\operatorname{NOD}(G,\alpha))$ is independent of the graph size.
-Observe that, also in this case, one class of inequalities is predominant, namely, (25).
-Notice also that some violated inequalities in (28) and (29) have only been detected for the densest graphs $p=0.9$.
+Observe that, also in this case, one class of inequalities is predominant, namely, (15b).
+Notice also that some violated inequalities in (16b) and (16c) have only been detected for the densest graphs $p=0.9$.
 Very good upper bounds are also achieved by $M_{+}(\operatorname{NOD}(G,\theta))$ when $p\geq 0.7$.
@@
 $n$ $p$ Time # iter Time # iter Time # iter Time # iter Time # iter Time
-200 0.1 7.24 1.6 26.64 3.2 48.24 0.2 8.74 0.2 8.64 0.2 8.76
+200 0.1 7.24 2.0 26.64 3.2 48.24 0.2 8.74 0.2 8.64 0.2 8.76
 0.2 4.29 1.0 9.10 2.0 13.96 0.0 4.34 0.0 4.34 0.0 4.34
@@
 0.4 1.74 0.2 2.65 2.0 6.51 0.0 1.82 0.2 2.13 1.0 3.77
-0.5 1.64 0.0 2.46 2.0 6.37 0.0 1.74 1.0 3.60 1.0 12.12
+0.5 1.64 0.0 2.46 2.0 6.37 0.0 1.74 1.0 3.60 1.8 12.12
 0.6 1.82 0.0 2.85 2.0 7.00 0.0 1.93 1.0 6.56 2.0 22.02
@@
 0.9 5.43 0.6 10.33 2.0 37.73 0.0 5.57 2.6 236.08 2.6 236.57
-250 0.1 9.62 1.0 28.92 3.0 56.76 0.0 9.68 0.0 9.68 0.0 9.68
+250 0.1 9.62 1.6 28.92 3.0 56.76 0.0 9.68 0.0 9.68 0.0 9.68
 0.2 5.15 1.0 11.53 2.0 18.05 0.0 5.25 0.0 5.25 0.0 5.25
@@
 0.8 3.78 0.0 7.42 1.0 10.76 0.0 4.31 2.2 97.62 3.6 197.18
-0.9 6.61 0.0 11.98 1.8 30.75 0.0 7.11 4.8 1250.18 5.0 1566.21
+0.9 6.61 0.0 11.98 1.8 30.75 0.0 7.11 4.8 1250.1 5.0 1566.21
 Table 4: CPU time and cutting plane’s iterations on random graphs.
@@
 $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\theta)$ $\nu(G,\alpha)$
-$n$ $p$ (12) (19) (25) (26) (28) (29) (25) (26) (28) (29)
+$n$ $p$ (11b) (13b) (15b) (15c) (16b) (16c) (15b) (15c) (16b) (16c)
 200 0.1 1018 3126 0 0 0 0 0 0 0 0
@@
 0.5 0 534 0 0 0 0 1822 178 0 0
-0.6 0 547 128 4 0 0 2828 172 0 0
-0.7 0 662 1772 60 0 0 3939 61 0 0
+0.6 0 547 128 4 0 0 $28$ 172 0 0
+0.7 0 662 1772 60 0 0 $39$ 61 0 0
 0.8 0 848 2138 62 0 0 3574 26 0 0
@@
 Table 7 shows that computing $\nu(G,\alpha)$ is about four times slower than computing $\theta^{+}(G)$ (including the time to build the LP relaxation), an acceptable price for such a valuable result.
-Table 8 also reveals the key contribution to closing the gap given by inequalities (25).
+Table 8 also reveals the key contribution to closing the gap given by inequalities (15b).
 Also in this experiment, the polynomial bound $\nu(G,\theta)$ improves on $\theta^{+}(G)$ in six cases by a meaningful amount (namely, DSJC125-5, MANN_a27, p_hat300-1, p_hat500-1, p_hat700-1).
@@
 Looking at Tables 7 and 8, one can observe that relevant improvements on $\theta^{+}(G)$ are achieved by adding several inequalities in a few iterations of the cutting plane algorithm.
-Similarly to what is observed for random graphs, inequalities (21) are only added in a few special graphs, namely the extremely sparse MANN graphs.
-Differently, classes of inequalities (25) and (26) are relevant to describe $M_{+}(\operatorname{NOD}(G,\theta))$, $M_{+}(\operatorname{NOD}(G,\alpha))$, while no violated members of classes (28) and (29) have ever been found.
+Similarly to what is observed for random graphs, inequalities (13d) are only added in a few special graphs, namely the extremely sparse MANN graphs.
+Differently, classes of inequalities (15b) and (15c) are relevant to describe $M_{+}(\operatorname{NOD}(G,\theta))$, $M_{+}(\operatorname{NOD}(G,\alpha))$, while no violated members of classes (16b) and (16c) have ever been found.
+Interestingly, our numerical results indicate that inequalities generated during the lifting step by multiplying by $x_{i}\geq 0$ are more frequently violated than those obtained by multiplying by $1-x_{i}\geq 0$.
+This trend is consistent across all reported relaxations, with only a few exceptions.
 Algorithm 2 has shown a nice numerical behavior even for the largest graphs in the DIMACS collection, and re-optimization is cost-effective (Table 7): its extra time is only paid when the cuts yield a significant improvement of the upper bound.
 In contrast, in the other cases, CPU times do not deviate significantly from those of $\theta^{+}(G)$.
-Overall, the computational analysis shows that the new SDP relaxations $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$, may be a viable option to compute upper bounds to $\alpha(G)$ which are remarkably stronger than $\theta(G)$ at a reasonable computational price.
-Interestingly, some experimental evidence complements the theoretical insights illustrated in Section 4.3, where $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ looks to be likely to achieve the tightest bounds than $M_{+}(\operatorname{NOD}(G,\alpha))$.
-7 Conclusions
-We have presented a way to improve the classical Lovász $\theta$ bound for the stability number of a graph by applying the Lovász-Schrijver $N_{+}(\cdot)$ lift-and-project operator to tailored LP relaxations.
-Unlike previous approaches, some of the latter may not be constructed in polynomial time.
-In the case of clique inequalities, we have to handle exponentially many inequalities (with an associated NP-hard separation problem), while building strong nodal inequalities requires solving the SSP on specific subgraphs.
-We show that these expedients, if properly handled, are indeed helpful in letting the resulting SDP relaxations improve the $\theta$ bound without excessive extra effort.
-This study reveals that the Lovász and Schrijver operators can be of practical interest besides being a powerful theoretical tool.
-Finally, an interesting research direction consists of finding classes of graphs where ${N_{+}(\operatorname{NOD}(G,\theta))}=\operatorname{STAB}(G)$, as for them the SSP would be polynomially solvable.
-Acknowledgements
-We wish to thank Adam Letchford and Stefano Gualandi for helpful discussions about the preliminary results of this study.
-Conflict of interest
-The authors declare that they have no conflict of interest.
+Overall, the computational analysis shows that the new SDP relaxations $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$, may be a viable option to compute upper bounds on $\alpha(G)$ which are remarkably stronger than $\theta(G)$ at a reasonable computational price.
+Interestingly, some experimental evidence complements the theoretical insights illustrated in Section 4.3, where $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ looks to be likely to achieve tighter bounds than $M_{+}(\operatorname{NOD}(G,\alpha))$.
 Graph $\theta^{+}(G)$ $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
-Name $d$(%) UB % gap UB % gap UB % gap UB % gap UB % gap UB % gap
+Name dd(%) UB % gap UB % gap UB % gap UB % gap UB % gap UB % gap
 brock200_1 25 27.20 29.508 27.20 29.505 27.12 29.119 27.20 29.508 27.20 29.508 27.20 29.508
@@
 Means - 30.832 - 30.528 - 30.248 - 30.832 - 30.489 - 27.772
-Table 6: Lift-and-project bounds on DIMACS instances. The column $d(\%)$ report the density of each graph.
+Table 6: Lift-and-project bounds on DIMACS instances. The column $d(\%)$ reports the density of each graph.
 Graph $\theta^{+}(G)$ $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
-Name $d$(%) Time # iter Time # iter Time # iter Time # iter Time # iter Time
+Name dd(%) Time # iter Time # iter Time # iter Time # iter Time # iter Time
 brock200_1 25 3.72 1 7.22 1 6.86 0 3.79 0 3.78 0 3.79
-brock200_2 50 1.63 0 2.51 1 4.02 0 1.73 1 3.79 1 14.22
+brock200_2 50 1.63 0 2.51 1 4.02 0 1.73 1 3.79 2 14.22
 brock200_3 39 2.07 1 4.27 1 4.44 0 2.15 1 3.92 1 4.47
@@
 brock800_4 35 19.86 0 50.82 0 27.46 0 23.98 0 24.00 2 76.29
-C125-9 10 3.46 2 15.68 3 32.44 1 7.10 1 7.44 1 7.55
-C250-9 10 10.45 1 32.52 2 42.68 0 10.52 0 10.51 0 10.52
+C125-9 10 3.46 2 15.68 4 32.44 1 7.10 1 7.44 1 7.55
+C250-9 10 10.45 2 32.52 2 42.68 0 10.52 0 10.51 0 10.52
 C500-9 10 25.12 1 57.41 1 53.55 0 25.52 0 25.50 0 25.50
-DSJC125.1 9 4.38 2 19.28 3 34.02 1 8.38 1 10.27 1 9.02
+DSJC125.1 9 4.38 2 19.28 4 34.02 1 8.38 1 10.27 1 9.02
 DSJC125.5 50 1.58 1 3.17 1 3.72 0 1.61 1 4.35 1 4.64
@@
 DSJC500-5 50 6.35 0 20.49 0 8.85 0 7.71 0 7.83 4 126.12
-MANN_a9 7 0.79 1 2.15 1 4.38 0 0.80 1 2.20 1 1.82
+MANN_a9 7 0.79 1 2.15 2 4.38 0 0.80 1 2.20 1 1.82
 MANN_a27 1 9.93 1 73.41 3 1992.72 0 10.00 1 57.22 1 60.32
@@
 p_hat300-1 76 9.07 0 12.62 1 22.93 0 9.49 2 134.81 3 261.52
-p_hat300-2 51 80.54 1 220.85 2 221.98 0 80.85 1 300.55 1 279.48
+p_hat300-2 51 80.54 2 220.85 2 221.98 0 80.85 1 300.55 1 279.48
 p_hat300-3 26 17.66 1 32.39 2 44.90 0 17.84 1 39.16 1 39.27
@@
 Graph $\lambda(G)$ $\mu(G,\mathcal{C})$ $\nu(G,\Gamma)$ $\nu(G,\theta)$ $\nu(G,\alpha)$
-Name (12) (19) (21) (26) (25) (26) (25) (26)
+Name (11b) (13b) (13d) (15c) (15b) (15c) (15b) (15c)
 brock200_1 10 471 0 0 0 0 0 0
@@
 p_hat700-2 1107 1549 0 0 16 286 718 282
-p_hat700-3 520 1414 0 0 0 47 1 187
+p_hat700-3 520 $14$ 0 0 0 47 1 187
 sanr200_0.7 3 286 0 0 0 0 0 0
@@
 Table 8: Number of added cutting planes on DIMACS instances.
+6.4 Sensitivity to the choice of $\mathcal{C}$
+To enhance the results of our computational analysis, we investigate whether refining the edge clique cover $\mathcal{C}$ obtained from Algorithm 1 can lead to a significant improvement in the bound $\mu(G,\mathcal{C})$, while also considering the associated computational overhead.
+To this end, we consider the edge clique cover $\mathcal{C_{\rm cut}}\in\mathcal{Q}$ proposed by Letchford et al.
+[40].
+The collection $\mathcal{C_{\rm cut}}$ is obtained by adding to $\mathcal{C}$ all cliques identified through a cutting plane algorithm that employs the separation heuristic of Marzi et al.
+[45].
+Since $\mathcal{C}\subseteq\mathcal{C_{\rm cut}}$, $\operatorname{QSTAB}(G,\mathcal{C_{\rm cut}})$ provides a tighter upper bound on $\alpha(G)$ than that of $\operatorname{QSTAB}(G,\mathcal{C})$}.
+Marzi et al.
+[45] and Letchford et al.
+[40] also show that this upper bound is very close to the bound provided by $\operatorname{QSTAB}(G)$}.
+However, this improvement may require a large number of clique inequalities.
+Tables 9 and 10 report the results of this sensitivity analysis on a subset of the previously tested DIMACS instances.
+We selected graphs of varying densities for which we expect that $\mathcal{C_{\rm cut}}$ has the potential to overturn the conclusion drawn from the previous experiment.
+As anticipated, the percentage gaps in Table 9 show an improvement of $\mu(G,\mathcal{C_{\rm cut}})$ over $\mu(G,\mathcal{C})$.
+This gain, however, generally comes at a substantial additional cost in terms of cutting-plane iterations and CPU time (up to a factor of 2), as reported in Table 10.
+Interestingly, $\mathcal{C_{\rm cut}}$ has a negligible effect on the bounds for dense graphs, where $\nu(G,\alpha)$ remains superior.
+Overall, this experiment suggests that the greedy edge clique cover provides a good trade-off between bound quality and computational time.
+A notable exception is keller4, where $\mu(G,\mathcal{C}_{\rm cut})$ attains a tight upper bound of $12.64$ on $\alpha(G)=11$, computed in a reasonable time.
+To the best of the authors’ knowledge, this represents the strongest bound reported in the literature; see, e.g., Locatelli [41], Giandomenico et al.
+[30] and references therein.
+7 Conclusions
+We have presented a way to improve the classical Lovász $\theta$ bound for the stability number of a graph by applying the Lovász-Schrijver $N_{+}(\cdot)$ lift-and-project operator to tailored LP relaxations.
+Unlike previous approaches, some of the latter may not be constructed in polynomial time.
+In the case of clique inequalities, we have to handle exponentially many inequalities (with an associated NP-hard separation problem), while building strong nodal inequalities requires solving the SSP on specific subgraphs.
+We show that these expedients, if properly handled, are indeed helpful in letting the resulting SDP relaxations improve the $\theta$ bound without excessive extra effort.
+This study reveals that the Lovász and Schrijver $N_{+}(\cdot)$ operator can be of practical interest besides being a powerful theoretical tools.
+Therefore, its application to other combinatorial optimization problems that admit strong SDP relaxations, such as the graph coloring problem, is a natural research line.
+From a theoretical standpoint, a few questions remain open.
+First, from Theorem 3 we have ${N_{+}(\operatorname{NOD}(G,\Gamma))}\subseteq\operatorname{TH}^{+}(G)$, whereas the numerical experiments suggest ${N_{+}(\operatorname{NOD}(G,\Gamma))}=\operatorname{TH}^{+}(G)$.
+However, it still remains an open question.
+Secondly, the experiments showed a strong predominance of violated constraints lifted by multiplying by $x_{i}\geq 0$ on those multiplied by $1-x_{i}\geq 0$.
+This deserves a further theoretical investigation.
+Finally, an interesting research direction consists of finding classes of graphs where ${N_{+}(\operatorname{NOD}(G,\theta))}=\operatorname{STAB}(G)$, as for them the SSP would be polynomially solvable.
+Acknowledgements
+We wish to thank Adam Letchford and Stefano Gualandi for helpful discussions about the preliminary results of this study.
+Conflict of interest
+The authors declare that they have no conflict of interest.
+Graph $\theta^{+}(G)$ $\mu(G,\mathcal{C_{\rm cut}})$ $\mu(G,\mathcal{C})$ $\nu(G,\alpha)$
+Name UB % gap UB % gap UB % gap UB % gap
+brock200_1 27.20 29.508 26.71 27.183 27.12 29.119 27.20 29.508
+brock200_2 14.13 17.758 13.98 16.510 14.12 17.674 14.02 16.795
+brock200_3 18.67 24.479 18.30 22.013 18.65 24.356 18.67 24.467
+brock200_4 21.12 24.242 20.78 22.246 21.10 24.101 21.12 24.242
+C125-9 37.55 10.431 35.87 5.487 36.23 6.566 37.55 10.430
+DSJC125.1 38.04 11.896 36.42 7.125 36.79 8.207 38.04 11.881
+DSJC125.5 11.40 14.021 11.04 10.359 11.36 13.598 11.35 13.531
+keller4 13.47 22.417 12.64 14.873 13.39 21.709 13.45 22.236
+p_hat300-1 10.02 25.253 9.81 22.622 10.02 25.194 8.58 7.288
+p_hat300-2 26.71 6.855 26.00 4.015 26.49 5.950 26.69 6.768
+p_hat300-3 40.70 13.056 39.74 10.384 40.44 12.321 40.70 13.053
+p_hat500-1 13.01 44.533 12.95 43.847 13.01 44.526 10.68 18.721
+p_hat500-2 38.56 48.306 37.94 45.934 38.38 47.597 38.54 48.225
+p_hat700-1 15.05 36.774 15.02 36.589 15.05 36.773 11.30 2.709
+p_hat700-2 48.44 10.091 47.89 8.831 48.24 9.647 48.41 10.023
+sanr200_0.7 23.63 31.296 23.22 28.976 23.59 31.048 23.63 31.296
+sanr200_0.9 48.90 16.439 47.49 13.061 48.03 14.356 48.90 16.440
+sanr400_0.5 20.18 55.217 20.17 55.142 20.18 55.216 19.10 46.951
+Table 9: Semidefinite bounds with improved edge clique covers.
+Graph $\mu(G,\mathcal{C_{\rm cut}})$ $\mu(G,\mathcal{C})$ $\nu(G,\alpha)$
+Name # iter Time # iter Time # iter Time
+brock200_1 2.0 10.60 1.0 6.31 0.0 3.34
+brock200_2 2.0 7.04 1.0 3.87 2.0 13.88
+brock200_3 2.0 7.03 1.0 4.18 1.0 4.26
+brock200_4 2.0 7.29 1.0 4.56 1.0 4.54
+C125-9 4.0 32.91 4.0 33.90 1.0 9.95
+DSJC125.1 4.0 32.81 4.0 33.64 1.0 9.22
+DSJC125.5 2.0 5.53 1.0 2.97 1.0 3.64
+keller4 4.0 29.15 1.0 11.57 1.0 15.01
+p_hat300-1 2.0 36.83 1.0 21.78 3.0 262.60
+p_hat300-2 3.0 311.08 2.0 168.35 1.0 232.25
+p_hat300-3 4.0 75.37 2.0 43.34 1.0 36.93
+p_hat500-1 1.0 37.54 1.0 40.53 5.0 854.09
+p_hat500-2 3.0 841.79 2.0 588.09 1.0 927.59
+p_hat700-1 1.0 76.46 1.0 78.54 7.0 6307.61
+p_hat700-2 3.0 1405.20 2.0 1332.51 1.0 1208.12
+sanr200_0.7 2.0 9.51 1.0 5.50 0.0 3.19
+sanr200_0.9 3.0 43.39 3.0 43.01 1.0 20.16
+sanr400_0.5 1.0 8.78 1.0 8.70 3.0 76.85
+Table 10: CPU time and cutting-plane iterations with improved edge clique covers.